### v1

# Comprehensive Benchmarking Environment for Worker Flexibility in Flexible Job Shop Scheduling Problems

Abstract
In Production Scheduling, the Flexible Job Shop Scheduling Problem (FJSSP) aims to optimize a sequence of operations and assign each to an eligible machine with varying processing times. To be even closer to real production environments, the integration of the workforce may be considered. Hence, each machine also requires a worker to be present to process an operation which additionally affects the processing times. The resulting problem is called Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W).
The FJSSP has been approached with different strategies considering various problem representations, including Mixed Integer Linear Programming (MILP), Constrained Programming (CP), and Simulation-based Optimization (SBO). In the latter area in particular, there exists a large number of specialized Evolutionary Algorithms (EA) like Particle Swarm Optimization (PSO) or Genetic Algorithms (GA). Yet, the solvers are often developed for single use cases only, and validated on a few selected test instances, let alone compared with results from state-of-the-art solvers using other problem representations. While suitable approaches do also exist, the design of the FJSSP-W instances is not standardized and the performance of the algorithms is hardly comparable.
This calls for a systematic benchmarking environment that provides a comprehensive set of FJSSP(-W) instances and supports targeted algorithm development. It will facilitate the comparison of algorithmic performance in the face of different problem characteristics. The present paper presents a collection of 402 commonly accepted FJSSP instances and proposes an approach to extend these with worker flexibility. The conversion of the FJSSP into an FJSSP-W has a significant impact on problem complexity. In addition, we present a detailed procedure for the evaluation of scheduling algorithms on these problem sets and provide suitable model representations for this purpose. This includes MILP formulations, CP models, and appropriate problem encodings for SBO solvers for both problem types. In addition, we provide complexity characteristics for all presented instances as well as baseline results of common commercial solvers to facilitate the validation of new algorithmic developments.
keywords:
Simulation-Based Optimization , Evolutionary Computation , Production Planning , Flexible Job Shop Scheduling , Worker Flexibility , Benchmarking
journal: Swarm and Evolutionary Computation
[…]
[inst1]organization=Josef Ressel Centre for Robust Decision Making
Research Centre Business Informatics
Vorarlberg University of Applied Sciences,addressline=
Hochschulstrasse 1, city=Dornbirn, postcode=6850, country=Austria
{highlights}
An alternative modeling approach for flexible job shop scheduling problems (FJSSP) as a mixed integer linear program (MILP) suitable to represent an exhaustive set of commonly accepted benchmark problems is proposed.
The paper presents a refined algorithmic approach to extend the FJSSP model with worker flexibility constraints (FJSSP-W). This way, we provide a portfolio of benchmark instances for an otherwise rather fragmented problem branch. A corresponding extension of the derived MILP representation for the introduced FJSSP-W is also delivered.
We provide a suitable problem encoding for a convenient simulation-based optimization (SBO) via meta-heuristics, e.g. Evolutionary Algorithm (EA) variants, to such problem classes and present an approach to effectively treat the worker flexibility constraints.
In addition, we develop adequate Constraint Programming (CP) representations for both problem classes (FJSSP and FJSSP-W) which facilitate the comparison with state-of-the-art solvers from this field of research.
Two problem classes in three representations (MILP, CP, and SBO) provide a comprehensive test environment for the comparison of solvers of different algorithmic paradigms. As a baseline, we compare the results of commercial state-of-the-art solvers (GUROBI, Hexaly, CPLEX (IBM), OR-Tools (Google)) and self-developed Genetic Algorithms (GA).
1 Introduction
Planning problems occur in a variety of application areas where the allocation and optimization of resources is critical to achieving specific goals. Consequently, planning problems are relevant in virtually all economic areas and the corresponding challenges are critical for ensuring sustainable economic growth, competitiveness, and efficient operations. In Supply Chain Management, for example, planning is essential for coordinating the flow of goods. In the Energy Sector, planning problems deal with the generation, distribution, and consumption of energy. Other application fields include Healthcare, Finance, Agriculture, and many more. Another important area is Manufacturing, i.e. the branch of industry that involves the production of finished products from raw materials. There, Production Scheduling (PS) deals with the allocation of resources such as machinery, labor and materials in order to achieve production targets efficiently. Such production goals include meeting customer requirements in terms of delivery times, minimizing idle time in certain production steps or responding to machine breakdowns. In this way, effective scheduling aims to promote sustainable development, drive innovation and ensure the resilience and competitiveness of manufacturing companies.
Production Scheduling problems involve determining the proper sequence of actions or decisions to attain (near) optimal performance on the shop floor. They take a variety of forms and are commonly approached by applying techniques from Artificial Intelligence (AI) or Operations Research (OR). A broad overview of the PS landscape is provided in Fig. 1, in which the problem branches on which this work is based are indicated by the blue boxes.
Figure 1: Localization of the problem classes under consideration in the area of Production Scheduling. Notice that other problem types exist and the illustration does by no means claim to be comprehensive.
Among such tasks, the Job Shop Scheduling Problem (JSSP) is a type of scheduling problem where multiple jobs (or tasks) share a common set of resources but have different processing requirements and routes through the production facility. Each job may require different machines or processes, and the order in which the jobs are processed can affect the overall efficiency of the system. JSSPs may turn out very complex due to the variability in job processing times, resource constraints, and the need to optimize the sequence of operations. Unlike the traditional JSSP, Flexible Job Shop Scheduling Problem (FJSSP) allows jobs to be processed on different machines. This provides more flexibility in resource allocation to those machines that possesses the required capabilities for a job. In this respect, the FJSSP represents an extension of the JSSP that considers a more dynamic and adaptable environment. The flexibility allows for more dynamic and adaptive scheduling, but it also introduces additional complexity to the scheduling problem, and optimization algorithms need to consider the various combinations of machine assignments to achieve an (near) optimal schedule. An additional layer of complexity to the scheduling process is added when worker-related dynamics are also considered within a FJSSP. Such problems are commonly referred to as Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W). There are many restrictions that may arise when considering worker flexibility in FJSSP, e.g., consecutive working hours, breaks or different qualifications of workers.
Depending on the modeling approach used and the optimization technique envisaged, JSSP and FJSSP model representations come in different forms. Common problem formulations include mathematical graphs, such as Resource-Task-Network (RTN) representations, permutation-based representation, or mathematical programming (MP) formulations in the form of objective functions and constraints. The latter usually come in the form of Mixed-Integer Linear Programming (MILP) problems and can be treated by various optimization techniques. Formulations as simulation-based optimization (SBO) problems are also frequently found in the literature. Instead of an analytical representation of the problem, the objectives and dependencies of a production line are implemented in the form of a black-box environment that allows the evaluation of individual production plans. Constraint Programming (CP) formulations are another way of expressing production planning problems. CP enables the integration of constraints and their solution mechanisms into a programming language. The problem is formulated in declarative form, while the solution process is secondary from the users’ point of view. This task is performed by a specialized constraint solver.
For SBO representations common approaches comprise meta-heuristic algorithms that rely on nature-inspired working principles, such as Genetic Algorithm (GA), Ant Colony Optimization (ACO) or Particle Swarm Optimization (PSO) variants. Hybrid approaches that combine multiple optimization techniques to exploit complementary strengths are also regularly considered. An example would be the combination of a GA with a Local Search method. Focusing on the exploration of the neighborhood of a given solution, Local Search (LS) solvers are used to refine solutions obtained from other methods and to find locally optimal solutions. In case of a MILP or CP formulation of the FJSSP, specialized commercial mathematical programming solvers of e.g. GUROBI and CPLEX (IBM) or constraint programming solvers like OR-Tools(Google), Hexaly or CP of CPLEX respectively, are widely used to find (near) optimal solutions. These solvers obtain state-of-the-art solutions for many FJSSP instances. Still, the choice of solver depends on factors such as the size and complexity of the FJSSP instance, the desired solution quality, and the available computational resources. Researchers and practitioners often experiment with multiple approaches to find the most effective solution for a specific FJSSP instance.
While the performance of most solvers is assessed on individual real-world problem instances and compared to selected solver variants only, some scheduling approaches are validated on rather small sets of test problem instances. In contrast to the real-world problems considered, the majority of FJSSP test sets must be considered fairly academic. Nevertheless, recent literature offers access to a number of standard FJSSP benchmark instances being spread across individual publications and modeled in different ways. Further, the benchmark problems usually vary in their structure and complexity. A unified benchmarking framework which allows dedicated algorithm performance comparisons on FJSSPs does not exist to the best of our knowledge. Moreover, there are hardly any comparable benchmark instances for FJSSP-W. As the modeling of FJSSP-W requires a careful consideration of both machine and worker-related constraints. The type of model affects the selection of solvers able to handle the increased complexity introduced by flexible scheduling policies. This introduces novel objectives like not only optimizing production efficiency but also taking into account the necessities of the workforce.
This paper proposes a systematic test environment of FJSSP and FJSSP-W instances to benchmark state-of-the-art solvers. For this purpose, an alternative MILP modeling approach for FJSSPs suitable for incorporation of a comprehensive set of 402 commonly accepted benchmark problems is introduced. Our MILP formulation allows for an improvement of the schedules found by the commercial solver GUROBI on some of the benchmark instances. The paper presents an algorithmic approach to extend the FJSSP model with worker flexibility constraints (FJSSP-W). This way, we provide a way to systematize the fragmented area of FJSSP-W instances. It provides an opportunity to create a portfolio of benchmark instances for the rather fragmented area of FJSSP-W instances. Further, suitable SBO problem encodings convenient for the application of meta-heuristics, e.g. Evolutionary Algorithm (EA) variants, as well as adequate CP representations are provided. As a result, a comprehensive test environment for the comparison of solvers from three different paradigms (MILP, CP, and SBO) on both problem classes (of 402 instances each) is established. Lastly, the paper describes a structured benchmark environment for the FJSSP(-W) which offers the possibility to choose among different problem characteristics, as well as the evaluation of a set of performance indicators (relative solution quality, wall clock time, etc.) and corresponding graphical visualizations. The test environment also provides characteristic attributes for each problem instance, such as the total number of operations, the flexibility of the machines or the problem-related duration variety, on the basis of which a targeted selection of sub-instances can be made. The functionalities are demonstrated by presenting baseline results of various state-of-the-art commercial solvers (GUROBI, Local Solver, CPLEX (IBM), OR-Tools (Google)) and a self-developed Genetic Algorithms (GA).
The remainder of this paper is structured as follows: In Section 2, a detailed description of the considered problem classes is provided. In addition, related work is discussed and available benchmarks as well as state-of-the-art solvers are presented. Section 3 deals with the formulation of the respective FJSSP representations. Next, the procedure for integration of worker flexibility is specified in Sec. 4. It realizes the transformation of an FJSSP instance into an FJSSP-W instance with just a few user-defined assumptions. Subsequently, the corresponding FJSSP-W formulations are specified. The detailed description of the benchmarking environment and its configuration options are the subject of Sec. 5. Afterwards, Section 6 presents a number of baseline results obtained by application of several state-of-the-art solvers. The contribution of this work and its impact on the field of research is discussed in summary in Sec. 7. Finally, Section 8 gives an outlook on future extensions of the presented benchmarking suite.
2 Problem Formulation & Related Work
The Flexible Job Shop Scheduling Problem (FJSSP) is the starting point for the following investigations. In this section, a formal definition of this problem is presented in Sec. 2.1 and current work on the FJSSP is outlined in Sec. 2.3. A large proportion of such research looks at very individual problems from real-world applications or only compares the developed solvers on individually selected benchmark instances. This complicates the performance comparison between the solvers significantly. In order to counteract this situation and to develop a structured benchmarking environment, Sec. 2.2 provides an overview of the available benchmark instances in the field of FJSSP and their characteristics.
2.1 Problem Formulation
The class of FJSSPs consists of NP-complete [1] combinatorial optimization problems that require to sequence a series of operations on a subset of available machines in such a way that a certain cost function $C$ is minimized. The machines are not necessarily homogeneous which allows for different processing times for each operation, depending on which machine it is assigned to.
In general formulation, the FJSSP introduces a set of $n$ jobs $\mathcal{J}=\{J_{1},\dots,J_{n}\}$ that need to be processed on a set of $m$ machines =$\mathcal{M}=\{M_{1},\dots,M_{m}\}$. Each of these jobs $J_{i}\>(i=1,\dots,n)$ consists of a predetermined sequence of $n_{i}$ operations. To this end, the notation $O_{i,j}$ refers to the $j$th operation within job $i$. Each of these operations $O_{i,j}$ is performed by one of the admissible machines in the set $\mathcal{M}_{i,j}\subset\mathcal{M}$. If operation $O_{i,j}$ is assigned to a specific machine, this machine is denoted by $M_{i,j}$. The total number of operations in a FJSSP can thus be calculated as $\texttt{N}=\sum_{i=1}^{n}n_{i}$. Accordingly, the ratio of the total amount of operations and the number of jobs gives the average number of operations per job $\frac{\texttt{N}}{n}$. The constant processing time of operation $O_{i,j}$ on machine $M_{k}\in\mathcal{M}_{i,j}\>(k=1,\dots,m)$ is given as $T_{i,j,k}$. The minimum processing duration among all possible operation-machine combinations of a single problem instance is referred to as ${T}_{\min}$. Conversely, the maximum processing duration is ${T}_{\max}$. Their difference defines the so-called duration span $\Delta T={T}_{\max}-{T}_{\min}$. In common scenarios, the general representation of an FJSSP assumes that all orders are known and all machines are operational at time 00.
In the context of FJSSPs, a large number of objective functions are conceivable with respect to the optimization of which the appropriate schedule must be composed. The performance criteria include minimizing the makespan, minimizing tardiness or maximizing resource utilization, to name but a few. While the makespan, i.e. the total time required for the execution of all jobs of a problem instance, is one of the most important objective functions in literature and it can be used as an illustrative example in the remainder of this paper, the objective function is to be represented as generally as possible. This allows the subsequent inclusion of various optimization objectives both in the problem modeling and in the performance analysis within the proposed benchmarking environment. Since every maximization task can be formulated equivalently as a minimization problem, w.l.o.g. it can be assumed that a given cost function is to be minimized in full satisfaction of the production constraints of the FJSSP formulated above.
Consequently, the cost function represents a mapping of a certain schedule representation $\mathbf{y}\in\mathcal{Y}$ to a particular cost function value $C(\mathbf{y})\in\mathds{R}$, i.e. $C:\mathcal{Y}\to\mathds{R}$. Hence, one is interested in the search space vector $\mathbf{y}$ that realizes a minimal $C(\mathbf{y})$ value. Note, that concrete examples of a parameter vector encoding for $\mathbf{y}$ is depending on the modeling approach and will be provided below.
As it is common in the field of meta-heuristic optimization, the expression fitness will be used synonymously for the quality of a discovered feasible solution $\mathbf{y}$, i.e. a feasible schedule, in the subsequent sections of this work. That is, solution candidates of higher fitness (or quality) correspond to solutions of lower cost function values. The quality of two admissible scheduling solutions can consequently be compared on the basis of their objective function values. Accordingly, lower cost function values $C(\mathbf{y})$ correspond to better production plans $\mathbf{y}\in\mathcal{Y}$. For a proper comparison of two specific schedules, an order relation $\prec_{c}$ is defined by
$\mathbf{y},\mathbf{z}\in\mathcal{Y}:\qquad\mathbf{y}\prec_{c}\mathbf{z}\>\ratio\Leftrightarrow\>C(\mathbf{y})<C(\mathbf{z}).$ (1)
Hence, the fitness of $\mathbf{y}$ would be rated higher than that of $\mathbf{z}$ in the case of Eq. (1). The corresponding smaller or equal relation $\preceq_{c}$ is defined analogously.
2.2 Existing FJSSP Benchmarks
This section discusses the existing benchmark instances for the Flexible Job Shop Scheduling Problem. While there are already a number of test problem collections available for the FJSSP, these are often not systematically used for the evaluation of the developed solvers. Thus, the number of algorithms that are benchmarked on a large number of different FJSSP instances is relatively small [2]. A majority of solvers are rather evaluated on problem subgroups originating from distinct sources, which means that the results of such studies are always subject to a latent problem selection bias. A large number of papers also only consider individual real-world problem representations. The algorithms developed in this way are highly specialized and are rarely compared carefully with established solvers. This has two potential implications: on the one hand, the advantages of rather good solvers may remain unclear if only demonstrated in a very specific setting and secondly, incremental approaches may be overvalued if they are not compared with state-of-the-art methods.
Taking into account the structure and characteristics of the instances in such benchmark sources, one observes that the proposed test collections differ significantly from each other. Some provide a large number of very similarly structured instances, others provide fewer instances but greater differences in terms of problem complexity.
These design differences between the benchmark collections complicates the selection of reasonable test instances. For example, it is difficult to decide which instances of what collection should be chosen to evaluate an algorithm on a group of rather complex FJSSP instances. Especially on such instances, performance advantages of meta-heuristics over commercial MILP solvers could be expected. Of course, the answer could be to test the solvers on all available instances. Yet, it can often be useful to consider specific problem groups separately in order to focus on the individual strengths of a solver. In cases where the exact optimum is already known, i.e. for instances of relatively low complexity, the computing time required to achieve the optimum would be of greater interest. On the other hand, improvements of the solution quality are desirable as long as the exact optimum is unknown.
In this work, all $402$ distinct FJSSP instances of the ten test problem collections in Table 1 are considered. The table also indicates the number of instances (# instances) contained in each collections as well as the average number of operations per job (# avg. ops per job). The latter is within the range of $3$ and $19.5$, indicating the differences between the proposed benchmarking sources. Apart from the benchmark sources in Table 1, other collections with close relationships must be named. These are the problem collections in [3] and [4]. The design in [4] is limited to problem instances of the classical Job Shop Scheduling Problem (JSSP) and thus renounces the flexibility to run any of the jobs on any machine. Conversely, the instances in [3] allow parallel processing of some operations within a job. Taking these instances into account would therefore make it necessary to extend the models. Due to the special properties listed, both benchmark sets range outside the scope of this paper and are thus excluded. Conclusively, to the best of our knowledge, the benchmark sources referred to in Table 1 represent an extensive set of established test instances matching the definition of the FJSSP in Sec. 2.1.
Table 1: An overview of ten FJSSP benchmark problem collections from literature.
No. Source name # instances # avg. ops per job Reference
1 BehnkeGeiger 60 5,0 [5]
2 Brandimarte 15 8,6 [6]
3 ChambersBarnes 21 11,7 [7]
4 DPpaulli 18 19,5 [8]
5 Fattahi 20 3,0 [9]
6 Kacem 4 3,2 [10]
7 Hurink Edata 66 8,9 [11]
8 Hurink Sdata 66 8,9 [11]
9 Hurink Rdata 66 8,9 [11]
10 Hurink Vdata 66 8,9 [11]
The usual specification of the production environments is carried out for each individual FJSSP instance in the form of a text file, which shows all the necessary information about the dependencies in the flexible job shop. An example is provided in Fig. 2 to facilitate the comprehension of the specifications. There, the 11th instance of the FJSSP collection from [7] is used to describe the problem encoding.
The first row of such a file commonly contains two (optionally three) numbers: the first value determines the number of jobs $n$ and the second one specifies the number of machines $m$ in the FJSSP instance. An optional 3rd value may indicate the average number of machines per operation. Although the latter is not necessary for the implementation of the problem instance. In the example above, the number of jobs is $n=10$, there are $m=10$ machine available, and each operations can on average be processed on $1$ machine.
Figure 2: Explanation of the standard FJSSP benchmark specifications on the basis of the 11th test instance from the collection in [7].
The remaining rows of the text file then represent the individual jobs $J_{i}$ $(i=1,...,n)$ of the problem instance. The first value in each row gives the number of operations $n_{i}$ to be executed in job $J_{i}$. The second number is the number of machines $\lvert\mathcal{M}_{i,1}\rvert$ available to process the first operation $O_{i,1}$. This is followed by a sequence of $\lvert\mathcal{M}_{i,j}\rvert$ pairs of numbers $(k,T_{i,1,k})$ indicating the index $k$ of the admissible machines $M_{k}$ as well as the processing times $T_{i,1,k}$ of operation $O_{i,1}$ on machine $M_{k}$. Next, the data for the second operation $O_{i2}$ are provided. In this way, all operations of all jobs are specified accordingly.
In order to facilitate a systematic classification of the 402 FJSSP test instances and consequently support an assessment of the evaluation results, the characteristic properties for each instance are given in the Appendix. In addition to the key properties such as the total number of jobs ($n$), machines ($m$), or operations(N) of an instance, seven other instance-specific statistics are provided that allow for a rough intuition of the individual benchmark instance complexity. These include the minimum ($T_{\min}$), maximum ($T_{\max}$) and average processing times ($\bar{T}$) measured over all possible machine-operation combinations, as well as the corresponding standard deviation $\sigma_{T}$. Moreover, the average number of operations per job $\frac{\texttt{N}}{n}$, the machine flexibility $\beta$, and the duration variety $dv$ are provided. The latter two indicators are used to additionally reflect the complexity of the FJSSP benchmark instances as they quantify the degree of flexibility and the differences in processing times an instance is exposed to.  [12]
The $\beta$-flexibility of a benchmark instance is determined as the ratio of the average amount of available machines per operation $m_{avg}$ and the total number of machines $m$ as
$\begin{split}\beta=\cfrac{m_{avg}}{m}\end{split}$ (2)
It indicates the degree of resilience in the FJSSP setting [13]. In case, all operations of all jobs can be processed by any machine, $\beta=1$ and the FJSSP is called fully flexible. Otherwise, i.e. $10<\beta<$ 1, it is only partially-flexible and, consequently, harder to solve. The duration variety $dv$ is defined as the quotient of the number of unique values of the processing times $T_{ijk}$ within a FJSSP instance $d_{unique}$ and the total number of machine assignment options across all operations, i.e.,
$\begin{split}dv=\frac{d_{unique}}{\sum_{i,j}\lvert\mathcal{M}_{i,j}\rvert}\end{split}$ (3)
Low $dv$ values indicate either very similar processing times across all available machines or a large amount of assignment options, respectively. This corresponds to many distinct scheduling solutions that have the same level of fitness.
Figure 3: Illustration of duration variety $dv$ and flexibility $\beta$ across all considered FJSSP instances.
Figure 3 displays scatter plots of two characteristic features of all benchmark instances. This provides an first overview of the diversity of the individual instances and reveals where similarities coincide. Details on the individual characteristics of the considered benchmark sources can be found in the Appendix. This rough insight into the benchmark properties alone supports the demand for a targeted filtering of instance subgroups to get a better insight into the algorithmic working principles of developed solvers.
2.3 Recent Work on Benchmarking Worker Flexibility in FJSSP
In order to retrieve benchmark instances including worker flexibility, the instances are usually specifically constructed for the given use case. In most cases, the constructed instances are based on subsets of commonly used benchmark instances [14][15][16]. In other cases, benchmark instances are randomly generated or constructed [16][17][18]. While the construction process is sometimes explained ([19][14]), they are either adapted for a specific use case or only use a subset of the commonly used benchmark problems. The algorithm used to enable the conversion of all FJSSP benchmark instances was developed similarly to the algorithm used by [19].
[18] tries to balance the makespan, maximum worker workload, and total workload of machines using an multi objective discrete teaching-learning based optimization approach. They use 30 newly constructed benchmark instances. [14] minimize makespan, labor cost and green production related factors using a non-linear integer programming model, as well as a non-dominated ensemble fitness ranking algorithm. The 31 benchmark instances used, are constructed based on a subset of the commonly used FJSSP benchmark instances. [15] use a simulation based approach to optimize the FJSSP with worker flexibility and uncertain processing times with the minimum makespan and minimum standard deviation of makespan as objective. 36 benchmark instances are used for evaluation, constructed based on a subset of the commonly used FJSSP instances. [16] compare a mixed-integer linear programming approach with NSGA-II in a distributed FJSSP with worker flexibility on a subset of the common benchmarks (50 instances) and randomly generated instances (8 instances). The objectives for their optimization are makespan minimization, balanced machine workload and workload of workers. [17] considers the JSSP with worker flexibility on newly generated benchmark instances, focusing on the maximum worker workload, minimizing the makespan and balancing the human energy expenditure. To evaluate their approach, 40 problem instances are generated.
3 Representations of the Flexible Job Shop Scheduling Problem
This section establishes the foundation for setting up the benchmarking environment by describing three different modeling approaches for the FJSSP. The first of these is a Mixed Integer Linear Programming (MILP) representation defined in Sec. 3.1. It allows exact methods of mathematical programming to be applied. Furthermore, an implementation as Constraint Programming (CP) model is provided in Sec. 3.2. Toolboxes for this type of problem representation are currently among the most sophisticated solving techniques for Production Scheduling tasks. Lastly, a convenient problem representation for Simulation-Based Optimization (SBO) algorithms, e.g. meta-heuristics, is presented in Sec. 3.3. It includes the definition of a precise schedule encoding and the design of the production system as an functional simulation environment.
3.1 Mixed-Integer Linear Programming Formulation
The Mixed Integer Linear Program (MILP) formulation is a proven model for the FJSSP to address these problem class using mathematical optimization methods, i.e. branch-and-bound or cutting-plane methods. In this section, the MILP is systematically formulated describing the decision variables, constraints and objective functions in detail.
Accordingly, the decision variables are defined in the following way:
Notation Description of decision variables Domain
$X_{i,j,i^{\prime},j^{\prime}}$ Indication of whether operation $O_{i,j}$ is processed after operation $O_{i^{\prime},j^{\prime}}$ on the same machine: $1$ if True, 00 else. $\mathds{Z}_{2}$
$Y_{i,j,k}$ The variable indicates whether operation $O_{i,j}$ is processed on machine $M_{k}$: $1$ if True, 00 else. $\mathds{Z}_{2}$
$C_{i,j}$ A variable representing the completion time of the $j$th operation $O_{i,j}$ in the $i$th job $J_{i}$ $\mathds{R}_{+}$
In this context, the notation $\mathds{Z}_{2}=\{0,1\}$ represents the ring of integers modulo $2$ indicating the binary character of a decision variable. Continuous variables, on the other hand, take on positive real numbered values in $\mathds{R}_{+}$. As a result, each schedule $\mathbf{y}\in\mathcal{Y}$ is represented by a feasible value assignment of these $(2+\texttt{N})\texttt{N}$ decision variables.
In addition, an index set $M_{i,j}$ is introduced that contains the indices $k$ of the permissible machines $M_{k}$ ($k=1,\dots,m$) for all operations $O_{i,j}$ ($j=1,\dots,n_{i}$) of all jobs $J_{i}$ ($i=1,\dots,n$):
$M_{i,j}=\{k:\text{$O_{i,j}$ can be performed on machine }M_{k}\}$ (4)
Next, the constraints of the FJSSP are described on the basis of these decision variables. To do so, the assignment, sequencing, and variable type constraints are considered independently. Afterwards, address a typical optimization objective associated constraint formulations.
3.1.1 General restrictions in FJSSP
Assignment constraints
It is imperative that the total number of $\texttt{N}=\sum_{i=1}^{n}n_{i}$ operations $O_{i,j}$ of an FJSSP instance be executed on exactly one of the machines available and admissible for the respective operation. These are specified by the indices in the set $M_{i,j}$.
[…]
By requiring that the sum of $Y_{i,j,k}$ over all admissible machine indices equals $1$, this set of N constraints guarantees that all operations $O_{i,j}$ are assigned to exactly one machine.
Sequencing constraints on the job level
The FJSSP demands that all operations of a single job need to be executed in a predefined sequence. For this type of restrictions, a novel formulation is proposed that involves the completion times $C_{i,j}$ of the operations within the FJSSP instance. Hence, one can require that the completion time $C_{i,1}$ of the first operation $O_{i,1}$ of job $i$ cannot be smaller than the corresponding processing time (see Eq. (6)). Also, the completion time of the $j$th operation $O_{i,j}$ of job $i$ cannot be smaller than the sum of the completion time of its predecessor operation $O_{i,j-1}$ in the same job $i$ and its own processing time (cf. Eq. (7)). Consequently, one receives another two sets of N constraints of the form
[…]
Sequencing constraints on machine level
Taking into account the order of operations on the machine level, it must be ensured that the completion times of two successor operations on the one and same machine are not unacceptably close to each other. For the proper formulation, an auxiliary large penalty parameter $L$ is introduced. This can, for example, be achieved by aggregating the maximal processing times of all operations over all machines
$L=\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n_{i}}\max\nolimits_{k=1,\dots,m}(T_{i,j,k}).$ (8)
Accordingly, the following two sets of constraints are obtained
$C_{i,j}\geq$ $C_{i^{\prime},j^{\prime}}+T_{i,j,k}-L\cdot(3-X_{i,j,i^{\prime},j^{\prime}}-Y_{i,j,k}-Y_{i^{\prime},j^{\prime},k})$ (9)
$C_{i^{\prime},j^{\prime}}\geq$ $C_{i,j}+T_{i^{\prime},j^{\prime},k}-L\cdot(X_{i,j,i^{\prime},j^{\prime}}+2-Y_{i,j,k}-Y_{i^{\prime},j^{\prime},k}).$ (10)
Both sets (9) and (10) are required to hold for operations $O_{i,j}$ with $1\leq i<n$ and $1\leq j\leq n_{i}$ as well as operations $O_{i^{\prime},j^{\prime}}$ with $i^{\prime}>i$ and $1\leq j^{\prime}\leq n_{i^{\prime}}$ on those machines $M_{k}$ that satisfy $k\in M_{i,j}\cap M_{i^{\prime},j^{\prime}}$.
The two sets of constraints (9) and (10) are trivially fulfilled as long as the bracketed factor multiplied by $L$ is nonzero. In case that operation $O_{i,j}$ and $O_{i^{\prime},j^{\prime}}$ are processed on the same machine $M_{k}$, i.e. $Y_{i,j,k}=1$ and $Y_{i^{\prime},j^{\prime},k}=1$, the sequencing between these two specific operations is controlled by the decision variable $X_{i,j,i^{\prime},j^{\prime}}$. If $X_{j,k,j^{\prime},k^{\prime}}=1$ holds, operation $O_{i,j}$ is executed after operation $O_{i^{\prime},j^{\prime}}$ on the same machine $M_{k}$. This implies that the first set of constraints defined by Eq. (9) is active. It ensures that the processing time of the successor operation $O_{i,j}$ does not start before the predecessor operation $O_{i^{\prime},j^{\prime}}$ is finished. Conversely, if $O_{i^{\prime},j^{\prime}}$ is executed after operation $O_{i,j}$ on the same machine, the decision variable is $X_{i,j,i^{\prime},j^{\prime}}=0$. That is, the second set of constraints (10) is active and guarantees non overlapping processing times of both operations, respectively.
Variable type constraints
This type of constraints restrict the assignment of decision variables to permissible values. They are specified by the definition of the decision variables above. The binary variables $X_{i,j,i^{\prime},j^{\prime}}$ and $X_{i,j,k}$ only accept the values 00 and $1$ in $\mathds{Z}_{2}$. The range of the continuous variables $C_{i,j}$ is restricted to positive real values in $\mathds{R}_{+}$. In summary
$X_{i,j,i^{\prime},j^{\prime}}\in\{0,1\},\qquad Y_{i,j,k}\in\{0,1\},\quad\text{and}\quad C_{i,j}>0.$ (11)
Without considering a specific objective function $C$ and potentially associated additional constraints, the restrictions of the FJSSP environment are now adequately described by the constraints in Eqs. (5), (6), (7), (9), (10), and (11).
3.1.2 Objective function and associated constraints
Makespan minimization
The total time required for the execution of all jobs in an FJSSP instance is referred to as the makespan of the FJSSP. Its minimization represents one of the most common objectives considered in this problem domain. The formal formulation of the makespan depends on the representation of the FJSSP and in this specific case coincides with the maximum value $C_{max}$ of the completion times $C_{i,j}$ of all operations $O_{i,j}$. That is, the objective becomes
$\min\big{(}C_{\max}\big{)}=\min\left(\max_{\tiny\begin{array}[]{c}1\leq i\leq n\\ 1\leq j\leq n_{i}\end{array}}C_{i,j}\right)$ (12)
When using the above definition of the makespan, it must be guaranteed that the completion time of the last operation $O_{i,n_{i}}$ in each job $J_{i}$ is not exceeding the makespan. Consequently, one gets an additional set of constraints N associated with this specific objective
$C_{max}\geq C_{i,n_{i}}\qquad\text{ for }1\leq i\leq n.$ (13)
3.2 Constraint Programming Formulation
The FJSSP can be easily formulated as a discrete finite constraint optimization problem(COP). The two COP solvers that we take into account in this comparison implement variable types, like interval variables and constraints like no-overlap or end-before-start, that are tailored to deal with scheduling problems.
In this comparison OR-Tools (ORT) an open-source solver developed by Google and CP-Optimizer (CPO) a constraint optimization solver provided as a part of the IBM ILOG CPLEX Optimization Studio are used. The problem formulation for both solvers uses the same subscripts, variables, and index sets as the MILP formulation (see 3.1).
3.2.1 Variables
The parameters representing the data are $T_{i,j,k}$ the processing time of operation $O_{i,j}$ on machine $k$. The variables are an interval variable $I(i,j)$ with starting time $start(i,j)$, ending time $end(i,j)$ and duration $duration(i,j)$ and an alternative interval variables $a\_I(i,j,k)$ for $k\in M_{i,j}$ to represent the feasible processing modes i.e. the selectable machines. For these intervals $a\_duration(i,j,k)=T_{i,j,k}$ holds. The ORT solver additionally uses a boolean variable $a\_presence(i,j,k)$ which indicates whether the alternative interval $a\_I(i,j,k)$ is chosen.
3.2.2 Objective
The objektive, i.e. minimizing the makespan, is formulated as
$\max_{1\leq i\leq n}(end\_of(I_{i,n_{i}}))\to\min$
3.2.3 Constraints
Assignment constraints
The assignment of an operation $O_{i,j}$ to a machine is done by linking the interval variable $I(i,j)$ to an alternative interval variable $a\_I(i,j,k)$.
This can be done by using the $alternative$ constraint of CPO. For $1\leq i\leq n$ and $1\leq j\leq n_{i}$ we add
$alternative\left(I(i,j),[a\_I(i,j,k):\mbox{ for }k\in M_{i,j}]\right).$
For the ORT solver the $exactly\_one$ constraint is used. For $1\leq i\leq n$ and $1\leq j\leq n_{i}$ we add
$exactly\_one$ $\left(a\_presence(i,j,k):\mbox{ for }k\in M_{i,j}\right)$
$start(i,j)$ $=a\_start(i,j,k)\mbox{ if }a\_presence(i,j,k)=True$
$duration(i,j)$ $=a\_duration(i,j,k)\mbox{ if }a\_presence(i,j,k)=True$
$end(i,j)$ $=a\_end(i,j,k)\mbox{ if }a\_presence(i,j,k)=True$
Sequencing constraints on job level
The sequencing constraints within jobs are formulated slightly different for the two solvers.
The CPO solver provides an $end\_before\_start$ constraint,
$end\_before\_start(I(i,j-1),I(i,j))\mbox{ for }1\leq i\leq n\mbox{ and }2\leq j\leq n_{i}$
and with the ORT solver starting and ending times are used
$start(i,j)\geq end(i,j-1)\mbox{ for }1\leq i\leq n\mbox{ and }2\leq j\leq n_{i}$
Sequencing constraints on machine level
The $no\_overlap$ constraint which is provided by both solvers makes it very easy to formulate the sequencing constraint within machines. For $1\leq k\leq m$ the following constraints are added:
$no\_overlap\left(a\_I(i,j,k):\mbox{ for }1\leq i\leq n,1\leq j\leq n_{i}\mbox{if }k\in M_{i,j}\right)$
3.3 Simulation-Based Optimization Attempt
Multiple problem encodings can be found in the literature, the most common one being [19, 12]. To encode the information using this system, the FJSSP is split into 2 vectors. The first vector contains the sequencing information while the second vector contains the machine assignments. The sequence vector is encoded using the job IDs rather than the specific operation IDs. Since all operations can only be scheduled after the previous operation in the job sequence is completed, the specific operation can be determined by the number of times the job ID has previously appeared in the sequence vector. The assignment vector contains the IDs selected for the operation in a fixed order. For example, for a problem with 2 jobs, job 1 consisting of 3 operations and job 2 consisting of 2 operations, the sequence vector must contain the values $\mathbf{\{1,1,1,2,2\}}$ in any order. Assuming there are 2 machines in the production environment and the instance is fully flexible (every operation can be performed on any machine), the assignment vector can contain any combination of $1$ and $2$, with a length corresponding to the amount of operations. The order of the values is fixed, starting from the first operation of the first job, moving to the last operation to the first job, then moving on to the first operation of the second job and moving through to the end to the last operation of the second job. One example for the described problem could look like Listing 1.
1s = [1, 2, 1, 1, 2]
2a = [2, 2, 1, 2, 1]
Listing 1: Example sequence and assignment vector
$\mathbf{s}=\begin{pmatrix}1&2&1&1&2\end{pmatrix}^{\top}\qquad\textrm{and}\qquad\mathbf{a}=\begin{pmatrix}2&2&1&2&1\end{pmatrix}^{\top}$ (14)
The resulting schedule is depicted in Fig. 4.
Figure 4: Example schedule produced by Listing 1.
Each schedule $\mathbf{y}=(\mathbf{s}|\mathbf{a})\in\mathcal{Y}\subset\mathds{N}^{\texttt{N}+\texttt{N}}$ is encoded by two distinct vectors $\mathbf{s}$, and $\mathbf{a}$. The sequence vector $\mathbf{s}\in\mathds{N}^{\texttt{N}}$ refers to the job sequence of the schedule and determines the order of the operations on their assigned machines. A feasible $\mathbf{s}$ contains each job index $i$ exactly $n_{i}$ times and the $j$th appearance of index $i$ corresponds to operation $O_{ij}$. Since the sequence of operations of a job is fixed, the job dependent sequence will always be feasible. The assignment vector $\mathbf{a}\in\mathds{N}^{\texttt{N}}$ contains the corresponding machine assignments. The order in $\mathbf{a}$ is fixed in the beginning and never changes to make sure each slot can only be assigned with available machines for every operation. This, together with the encoding chosen for $\mathbf{s}$ ensures that all solutions produced either randomly or through recombination and mutation will always be feasible. To evaluate the solution, the schedule is constructed according to the instructions in $\mathbf{s}$ and $\mathbf{a}$. To construct the schedule, the operations are inserted according to the sequence in $\mathbf{s}$. The start time $t_{ij}$ is determined based on the current last operation ending time on the assigned machine $t_{a_{ij}}$ and the last ending time in the job sequence of the operation $t_{prev}$ (Eq. (15)).
$t_{prev}=\begin{cases}t_{i,j-1}+d_{i,j-1}&\text{if $j>0$}\\ 0&\text{if $j=0$}\end{cases}$ (15)
$t_{ij}$ is then determined by Eq. (16).
$t_{i,j}=\max(t_{prev},t_{a_{i,j}}).$ (16)
$t_{a_{,ij}}$ is replaced (Eq. (17)) and the next operation is inserted into the schedule.
$t_{a_{i,j}}=t_{i,j}+d_{i,j}.$ (17)
After all starting times have been determined, the solutions fitness $C_{max}$ is determined by Eq. (18).
$C_{\max}=\max_{\tiny\begin{array}[]{c}1\leq i\leq n\\ 1\leq j\leq n_{i}\end{array}}(t_{i,j}+d_{i,j})$ (18)
4 Expansion of the FJSSP for Worker Flexibility
The Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W) is an extension to the FJSSP which adds worker flexibility. This means that for every operation, an eligible worker also needs to be assigned to operate the machine to which the operation is assigned. The selected machine-worker combination further influences the processing time of the operation on the machine. In addition to the FJSSP, the FJSSP-W therefore introduces a set of workers $\mathcal{W}=\{W_{1},\dots,W_{o}\}$. Each operation $O_{i,j}$ needs to be assigned to a permissible machine $M_{k}\in\mathcal{M}_{i,j}\subset\mathcal{M}$ and a worker $W_{s}\in\mathcal{W}_{i,j,k}\subset\mathcal{W}$ the set of permissible workers to process the operation $O_{i,j}$ on the assigned machine $M_{k}$. Similar to the FJSSP, the FJSSP-W assumes that all operations, machines, and workers are operational and available at time 0. The objective for the FJSSP-W is to find those operation-machine-worker combination that minimize the makespan $C$. As opposed to the FJSSP, the flexibility $\beta$ of a benchmark instance needs to take more into account than only the machines. All possible assignment combination for every operation need to be taken into account. This is the same as for the FJSSP, but includes the possible worker assignments for each operation-machine combination as well. The average amount of assignment options $\omega_{avg}$ is determined by Eq. (20).
$\omega_{all}=\sum_{i=1}^{\texttt{N}}(|\mathcal{M}_{i}|+\sum_{j=1}^{|\mathcal{M}_{i}|}|\mathcal{W}_{\mathcal{M}_{i},j}|)$ (19)
$\omega_{avg}=\frac{1}{\texttt{N}}\sum_{i=1}^{n}\sum_{j=1}^{n_{i}}\sum_{k=1}^{m}|\mathcal{W}_{ijk}|\quad$ (20)
Which leads to $\beta$ defined as Eq. (21), with $\omega_{uni}$ being the overall amount of unique machine-worker combinations available for the benchmark instance.
$\beta=\cfrac{\omega_{avg}}{\omega_{uni}}$ (21)
In the same vein, the definition for $dv$ for FJSSP-W needs to be rewritten as Eq. (22)
$\begin{split}dv=\cfrac{d_{unique}}{\sum_{i=1}^{\texttt{N}}(|\mathcal{M}_{i}|+\sum_{j=1}^{|\mathcal{M}_{i}|}|\mathcal{W}_{\mathcal{M}_{i},j}|)}\end{split}$ (22)
4.1 Design Proposal of Worker Flexibility Integration
The FJSSP-W considers that operations in the production process need workers to be present. Therefore, workers also need to be assigned to the different operations. In the FJSSP-W, it is considered that not all workers will work at the same speed, and not all workers are capable of performing every operation with ever (or any) of the available machines for the operation. To receive a set of benchmark problems to test and develop appropriate approaches, the given FJSSP benchmark instances are extended to contain the additional required information about the possible worker assignments. To that end, we provide an algorithm similar to [20] , which is capable of extending any FJSSP instance in the format of the gathered benchmark instances (Alg. 1). The parameters used for the benchmark generation are the amount of workers (default: $1.5*m$), the upper and lower bound range for the durations (default: $ub=1.1$, $lb=0.9$ of the original duration) to have a random fluctuation in the durations required by the workers of ±$\pm 10\%$, and optionally a random seed for reproducibility.
After the initialization of the required parameters, in line 1 an empty list for saving the result benchmark is created. In line 2, the original benchmark file, located at $path$, is read and saved into a list, separated into the rows of the file. The base information about the benchmarks are added to the result, as a list representing a new row in the output, in line 3, keeping the same format as the original files, with the added information about the amount of workers available for the problem instance. Starting from line 4 and ranging to line 25, the rest of the rows of the original benchmark file are processed one by one. In line 5, a variable to keep track of the reading position in the current row ($idx$) is initialized, as well as an empty list, which is added to the output as a new row. The amount of operations needed for the job represented by the row is added to the new row in line 6. From line 7 to line 24, the current row from the original benchmark, according to the original format of the problem instances, is processed. To do so, $idx$ is raised by $1$ in row 8 to address the next operation that needs to be parsed. In line 9, the amount of machines available for the operation is read, and added to the output row in line 10. From line 11 to line 23, the different machine assignments and their respective durations for the current operation are parsed. To achieve that, line 12 raises $idx$ by $1$ again to advance the reading position of the input current row. Since $idx$ is now referencing to a machine ID in the original benchmark, this possible machine assignment is also added to the output row in line 13. In line 14, the amount of workers which are going to be able to perform the required operation on this specific machine is chosen uniformly from the interval [$1$, $n_{workers}$]. The amount of workers available for this machine assignment ($workers$) is added to the output row in line 15. In line 16, the available workers for the assignment are selected randomly, using a function named $Sample$, using the parameters $n_{workers}$ and $workers$ to select the IDs of the workers in the interval [0, $n_{workers}$). Note that generally the assigned IDs of machines and workers are converted to 0-based indexing for easier implementation as opposed to the original 1-based indexing. The subset of workers is stored in the list $\mathbf{selected}$. To advance the reading position again, $idx$ is raised by $1$ in line 17. $idx$ now points at the original duration of the machine assignment for the given operation, which is stored in $d_{orig}$ in line 18. Starting from line 19 and ending in line 22, the current operation-machine assignments are finalized by choosing the durations for the different worker assignments. To do this, the duration for each worker $d_{work}$ is uniformly chosen from the interval [$d_{orig}*lb$, $d_{orig}*ub$] and rounded to the closest integer in line 20. Line 21 adds the worker and the duration for the operation-machine-worker combination to the output row in $out$.
Algorithm 1 Adding Worker Flexibility to common FJSSP Benchmark instances
1:$n_{workers}>0$, $lb\geq 0$, $ub\geq 0$, $path$
2:$\mathbf{out}\leftarrow\{\}$
3:$\mathbf{rows}\leftarrow ReadFile(path)$
[…]
5:for $i\leftarrow 1$ to $||\mathbf{rows}||$ do
6:     $\mathbf{out}_{i}\leftarrow\{\}$, $idx\leftarrow 0$
7:     $\mathbf{out}_{i}\cup\mathbf{rows}_{i,idx}$
8:     for $j\leftarrow 1$ to $||\mathbf{rows}_{i}||$ do
9:         $idx\leftarrow idx+1$
10:         $m\leftarrow\mathbf{rows}_{i,idx}$
11:         $\mathbf{out}_{i}\cup\{m\}$
12:         for $k\leftarrow 1$ to $m$ do
13:              $idx\leftarrow idx+1$
14:              $\mathbf{out}_{i}\cup{\mathbf{rows}_{i,idx}}$
15:              $workers\leftarrow Uniform(1,n_{workers})$
16:              $\mathbf{out}_{i}\cup\{workers\}$
17:              $\mathbf{selected}\leftarrow Sample(n_{workers},workers$)
18:              $idx\leftarrow idx+1$
19:              $d_{orig}\leftarrow\mathbf{rows}_{i,idx}$
20:              for $l\leftarrow 1$ to $workers$ do
21:                  $d_{work}\leftarrow Round(Uniform(d_{orig}*lb,d_{orig}*ub))$
22:                  $\mathbf{out}_{i}\cup\{selected_{l},d_{work}\}$
23:              end for
24:         end for
25:     end for
26:end for
After this routine is finished, the benchmark is extended with worker flexibility and kept similar to the original format. The original format can be seen in an arbitrarily constructed example in Listing 2.
12 3
23 2 0 10 2 15 3 0 10 1 5 2 10 2 0 10 2 15
32 1 1 15 2 1 10 2 15
Listing 2: Original Benchmark example
Row 1 of Listing 2 contains the information that the problem instance consists of 2 jobs and 3 available machines. In row 2 it is specified that Job 1 consists of 3 operations. Operation 1 of Job 1 can be processed on 2 machines (Machine 0 with processing time 10 and Machine 1 with processing time 15). Operation 2 of Job 1 can be processed on 3 machines. Either Machine 0 with processing time 10, Machine 1 with processing time 5 or Machine 2 with processing time 10. Operation 3 of Job 1 has 2 machines available (Machine 0 with processing time 10 and Machine 2 with processing time 15). Job 2 requires 2 operations. Operation 1 of Job 2 can be processed on only 1 machine (Machine 1 with processing time 15), while Operation 2 of Job 2 can be processed on 2 machines (Machine 1 with processing time 10 and Machine 2 with processing time 15).
In the new format of the benchmarks, line 1 contains the general information about the instance (amount of jobs, amount of machines, amount of workers), followed by a row by row description of every job that needs to be scheduled. Every row starts with the number of operations required for the job to be finished. Following this first row entry, every operation starts with the amount of machines required, followed by the entries for every machine. Every entry for the machine starts with the machine id, followed by the amount of workers available to process the operation on this specific machine. After that, for every worker available on the machine, an entry consisting of the worker id, followed by the duration of the operation-machine-worker combination is listed in the row.
An arbitrarily constructed example for this format can be seen in Listing 3. REPLACE:
1 2 3 4
2 3 2 0 2 1 10 3 5 2 1 0 15 3 0 1 2 10 1 2 0 10 3 5 2 2 2 5 3 15 2 0 2 2 15 3 10 2 2 0 15 1 5
3 2 1 1 3 1 10 2 5 3 10 2 1 1 2 10 2 2 0 10 3 15
Listing 3: Benchmark with worker flexibility example
This example means that, according to row 1, there are 2 jobs, 3 machines and 4 workers in this production environment. Job 1 (row 2) consists of 3 operations. Operation 1 can be processed on 2 machines. The first of these machines is Machine 0, which has 2 workers available to perform this operation. Worker 1 with a processing time 10 and Worker 3 with a processing time of 5. The second machine to process Operation 1 of Job 1 is Machine 2. Machine 2 has 1 worker available (Worker 0) with a processing time 15. Operation 2 of Job 1 has 3 machines available. Machine 0, with 1 worker (Worker 2) with processing time 10, Machine 1 with 2 workers (Worker 0 with processing time 10 and Worker 3 with processing time 5), and Machine 2 with 2 workers (Worker 2 with processing time 5 and Worker 3 with processing time 15). Operation 3 of Job 1 can be processed on 2 machines. Machine 0, which has 2 possible worker assignments (Worker 2 with processing time 15 and Worker 3 with processing time 10), and Machine 2. Machine 2 also has 2 possible workers (Worker 0 with processing time 15 and Worker 1 with processing time 5). Job 2 (row 3), requires 2 operations to be processed. Operation 1 of Job 2 has 1 possible machine (Machine 1) with 3 possible workers. Worker 1 with processing time 10, Worker 2 with processing time 5 and Worker 3 with processing time 10. Operation 2 of Job 2 can be assigned to 2 machines. Either Machine 1, which has 1 possible worker (Worker 2 with processing time 10), or Machine 2, with the possibility to assign the operation to 2 different workers (Worker 0 with processing time 10 or Worker 3 with processing time 15).
4.2 Mixed-Integer Linear Programming Formulation
An obvious generalization of the MILP formulation for the FJSSP (3.1) is used. The worker flexibility is modeled by an additional set of constraints that follows the same ideas as the sequencing constraints on machine level in 3.1. The following description explains in detail only the additional notations and constraints of the MILP formulation for the FJSSP-W.
4.2.1 Indices and Variables
To denote the workers the subscripts $s,s^{\prime}$ with $1\leq s,s^{\prime}\leq w$ and $w$ the number of workers are used. The processing time of operation $O_{i,j}$ on machine $M_{k}$ also depends on worker $W_{s}$ and is denoted by $T_{i,j,k,s}$
The decision variables $Y_{i,j,k}$ in 3.1 have to be changed to $Y_{i,j,k,s}$ which are 1, if operation $O_{i,j}$ is processed on machine $M_{k}$ by worker $W_{s}$ and 0 else. Furthermore an additional type of decision variables is necessary: $U_{i,j,i^{\prime},j^{\prime}}$, which is set to 1 if Operation $O_{i,j}$ is processed after operation $O_{i^{\prime},j^{\prime}}$ by the same worker. To describe the admissible workers for processing operation $O_{i,j}$ on machine $M_{k}$ the index set
$W_{i,j,k}=\{s:\text{ such that $O_{i,j}$ can be processed on $M_{k}$ by worker$W_{s}$}\}$
is used.
4.2.2 Constraints
Assignment constraints
These constraints guarantee that all operations $O_{i,j}$ are assigned to exactly one machine and one worker.
$\sum\nolimits_{k\in M_{i,j}}\sum\nolimits_{s\in W_{i,j,k}}Y_{i,j,k,s}=1\text{ for }1\leq i\leq n,1\leq j\leq n_{i}$
Sequencing constraints on job level
$C_{i,1}\geq\sum\nolimits_{k\in M_{i,1}}\sum\nolimits_{s\in W_{i,1,k}}Y_{i,1,k,s}T_{i,1,k,s}\text{ for }1\leq i\leq n$
$C_{i,j}\geq C_{i,j-1}+\sum\nolimits_{k\in M_{i,j}}\sum\nolimits_{s\in W_{i,j,k}}Y_{i,j,k,s}T_{i,j,k,s}\text{ for }1\leq i\leq n,1<j\leq n_{i}$
Sequencing constraints on machine level
$C_{i,j}\geq C_{i^{\prime},j^{\prime}}+T_{i,j,k,s}-L(3-X_{i,j,i^{\prime},j^{\prime}}-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k,s^{\prime}})$
$C_{i^{\prime},j^{\prime}}\geq C_{i,j}+T_{i^{\prime},j^{\prime},k,s^{\prime}}-L(X_{i,j,i^{\prime},j^{\prime}}+2-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k,s^{\prime}})$
Both constraints have to hold for operations $O_{i,j}$ with $1\leq i<n$ and $1\leq j\leq n_{i}$ and operation $O_{i^{\prime},j^{\prime}}$ with $i^{\prime}>i$ and $1\leq j^{\prime}\leq n_{i^{\prime}}$, for machines $M_{k}$ with $k\in M_{i,j}\cap M_{i^{\prime},j^{\prime}}$ and for workers $W_{s}$ ($s\in W_{i,j,k}$) and $W_{s^{\prime}}$ ($s^{\prime}\in W_{i^{\prime},j^{\prime},k}$).
Sequencing constraints on worker level
$C_{i,j}\geq C_{i^{\prime},j^{\prime}}+T_{i,j,k,s}-L(3-U_{i,j,i^{\prime},j^{\prime}}-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k^{\prime},s})$
$C_{i^{\prime},j^{\prime}}\geq C_{i,j}+T_{i^{\prime},j^{\prime},k^{\prime},s}-L(U_{i,j,i^{\prime},j^{\prime}}+2-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k^{\prime},s})$
As in 4.2.2 these two sets of constraints have to hold for operations $O_{i,j}$ with $1\leq i<n$ and $1\leq j\leq n_{i}$ and operation $O_{i^{\prime},j^{\prime}}$ with $i^{\prime}>i$ and $1\leq j^{\prime}\leq n_{i^{\prime}}$. But contrary to 4.2.2 they have to hold for machines $M_{k}$ ($k\in M_{i,j}$) and $M_{k^{\prime}}$ ($k^{\prime}\in M_{i^{\prime},j^{\prime}}$) and for workers $W_{s}$ with $s\in W_{i,j,k}\cap W_{i^{\prime},j^{\prime},k^{\prime}}$. For the large number $L$, one can use
$L=\sum\nolimits_{i=1}^{n}\sum\nolimits_{j=1}^{n_{i}}\max\nolimits_{1\leq k\leq m;1\leq s\leq w}(T_{i,j,k,s})$
The idea of these constraints is the same as in Eq. (9) and (10) explained there in detail. Finally makespan and variable type constraint are the same as in 3.1.
4.3 Constraint Programming Formulation
The same as for the MILP formulation holds also for the CP formulation: The FJSSP-W is an obvious generalization of the FJSSP. In the following the formulation for CPO solver is presented. The necessary changes for using the ORT solver are obvious when the two different formulations for the FJSSP are kept in mind. Again, each operation is represented by an interval variable $I(i,j)$ with starting time $start(i,j)$, ending time $end(i,j)$ and duration $duration(i,j)$. For the FJSSP-W an alternative interval variable $a\_I(i,j,k,s)$ has to be defined for each admissible machine-worker combination.
The assignment of an operation $O_{j,k}$ to exactly one machine-worker combination is done by linking the interval variable $I(j,k)$ to an alternative interval variable $a\_I(i,j,k,s)$. This can be done by the $alternative$ constraints
$alternative\left(I(i,j),[a\_I(i,j,k,s):\mbox{ for }k\in M_{i,j},s\in W_{i,j,k}]\right)$
for $1\leq i\leq n$ and $1\leq j\leq n_{i}$.
The sequencing constraints on job level are the same as in 3.2:
$end\_before\_start(I(i,j-1),I(i,j))\mbox{ for }1\leq i\leq n\mbox{ and }2\leq j\leq n_{i}$
and the sequencing constraints on machine level are very similar to 3.2. For $1\leq k\leq m$ the following constraints are added:
$no\_overlap\left(a\_I(i,j,k,s):\mbox{ for }1\leq i\leq n,1\leq j\leq n_{i}\mbox{ if }k\in M_{i,j}\mbox{ and }s\in W_{i,j,k}\right)$
Additionally the sequencing on worker level has to be guaranteed by using another set of $no\_overlap$ constraints. For $1\leq s\leq w$:
$no\_overlap\left(a\_I(i,j,k,s):\mbox{ for }1\leq i\leq n,1\leq j\leq n_{i},k\in M_{i,j}\mbox{ if }s\in W_{i,j,k}\right)$
4.4 Simulation-Based Optimization Attempt
The extension to the encoding for simulation-based approaches stays the same for sequence and machine assignments as in section 3.3. To add the worker flexibility, an additional vector is added. The vector has the same properties as the machine assignment vector, in that it is in the same fixed order and contains the IDs of the workers assigned to the operation-machine combination chosen to do the task. Extending the example used in section 3.3 with 3 available workers for the system, a problem instance could look like Listing 4.
1s = [1, 2, 1, 1, 2]
2a = [2, 2, 1, 2, 1]
3w = [3, 1, 1, 2, 3]
Listing 4: Example sequence, assignment and worker vector
The resulting schedule looks as depicted in Fig. 5.
Figure 5: Example schedule produced by Listing 1.
5 Description of the Benchmarking Environment
5.1 Implementation details
The benchmarking environment is implemented in Python [21]. Several APIs are offered by the environment, including the evaluation of FJSSP and FJSSP-W solutions and their comparison to the known results.
Additionally, the benchmarking environment contains the best known values for the FJSSP and FJSSP-W instances created with the same parameters than the instances used in this paper. The algorithm which turns FJSSP instances into FJSSP-W instances is also provided.
5.2 Performance evaluation
For comparisons to other solvers the benchmark environment offers different visualizations of the results over all used benchmark instances. The data of the evaluated solvers are included and best known results can be updated over time.
(a)Refer to caption (b)Refer to caption
Figure 6: An example plot using example data displaying the portion of instances that lie within a certain gap from the best found solution. A gap of 1.0 corresponds to 100% distance. In (a), all of the data is displayed, while (b) shows an example with a more zoomed in perspective on the same data (limit on the x-axis as parameter to the visualizer code)
The example plot shown in Fig. 6(a) compares the relative deviation (gap) $\delta_{rel}$ (Eq. (23)) of the best solution of a solver $C_{fb}=C(\mathbf{y}_{fb})$ for each algorithm from the best-known result $C_{best}$. On the y-axis, the plot shows the portion of instances where the solutions were within the given gap to $C_{best}$.
$\begin{split}\delta_{rel}=\frac{C_{fb}-C_{best}}{C_{best}}.\end{split}$ (23)
Fig. 6(b) shows the same (on example data) with a zoomed in perspective (limit on the x-axis).
(a)Refer to caption (b)Refer to caption
Figure 7: Based on example data, the portion of problem instances per solver that were optimized within 10% of the best found result. (a) shows the results in relation to the best solution of all solvers within, while (b) shows the time needed for the instances to reach the threshold in relation to their own best solution.
The plots shown in Fig. 7 show the portion of instances which reached the given criteria in $n$ seconds. The example plots use the fitness as example, but the same plots can also be done for other criteria, e.g. the lower bounds found by a solver. The maximal distance to the target value can be given to the visualizer as a parameter. In the case of Fig. 7(a), the plot shows the portion of instances capable of getting within a range of no more than 10% distance from the overall best known result $C_{best}$ within the given time frames. Fig. 7(b) however shows how long it took each solver to get within the range of 10% of it’s own best solution $C_{fb}$ for which portion of the test instances. Note that while for Fig. 6(a) all solvers reach 100% eventually, granted that all instances included in the comparison have been solved by each solver, in Fig. 7(a) there may be many instances that never reached the threshold. The visualizer is implemented in a way that it can handle missing data for the case of Fig. 6(a). The corresponding line for the solver with the missing results would not reach 100%, which is the only way this can happen for these plots (without a limit imposed on the x-axis).
6 Baseline Experiments
6.1 Experimental Settings
For the FJSSP-W, the solvers were tasked with solving all 402 benchmarks. The hardware used for the experiments is a Intel Core i7-6700 (3.40GHz) CPU with 16 GB RAM. All solvers were given a maximum of 1200 seconds (20 minutes) for each benchmark instance. The benchmark instances, including all results are available online [22].
6.2 Applied Solvers
All used solvers, including the self implemented GA, were not manually tuned or configured to accommodate any of the benchmark instances and were run in their default settings. However, some of the solvers do automated pre-processing and parameterizing on their own, e.g. Gurobi or the self implemented GA, which determines some of its parameters automatically based on the benchmark characteristics.
Hexaly
Hexaly is a new kind of global optimization solver. It offers nonlinear and set-oriented modeling APIs. Relying on innovative proprietary algorithms, it is faster and more scalable than Mixed-Integer Programming, Constraint Programming, and Nonlinear Programming solvers. [23]
OR-Tools
OR-Tools is an open-source software suite for optimization [24] developed by Google. It offers several ways to solve constraint programming and linear programming problems. The solver used for the results in this paper is the CP-SAT solver, which is a constraint programming solver that uses SAT (satisfiability) methods.
ILOG CPLEX
CPLEX by IBM is an optimization software package used to solve several optimization problem types, including integer programming, linear programming, quadratic programming problems, and constraint programming problems. CPLEX is used for both its constraint programming capabilities (CPLEX-CP) and its MILP solver capabilities (CPLEX-LP). [25]
Gurobi
Gurobi is a commercial mathematical solver commonly used to solve MILP problems and is developed by Gurobi Optimization LLC. [26]
Genetic Algorithm
To solve the FJSSP-W, an extended version of the GA implemented in [12] is used. Additionally, the implementation is done in the programming language C# rather than Python. To evaluate the GA properly, all instance are repeated 18 times to account for the randomness of the GA. The 18 repetitions are a result of the employed hardware being able to handle 6 instances of the GA running in parallel processes. With a maximum running time of 20 minutes, all repetitions for all instances were expected to take 402 hours (16.75 days). However, this runtime was reduced by preemptively stopping the GA once it reached the optimal result. However, other than the CP and MILP solvers, the GA has no built-in method to determine if a found solution is optimal. Therefore, the GA was run as the last solver and relied on results reported by the other solvers as optimal results, if any optimal result was found for a problem instance. This helps illustrate which ranges of the different characteristics are easiest to handle. For example, it can be observed that the difficulty of a problem increases with a higher number of operations, as well as with a lower duration variety in the processing times.
6.3 Baseline Results
Figure 8: CDF plot showing selected characteristics of all problem instances
The tables for each solver show the number of instances in which the solver reached the best result, the average characteristics of the instances and the respective standard deviations. The rows in the tables separate the instances in which the solver was the only solver to reach the best result (Single Best), the instances in which at least one other solver also reached the best result (Shared Best), an overview of all instances for which the solver found the best result (Combined), which includes the 2 prior rows, and lastly the instances for which the solver reached the optimal solution (Optimal). Note that since the GA has no way to confirm if a solution is an optimum, the best results for each instances were collected. For each solver the instance counts if it reached the known optimum even if it did not identify it as the optimum, as long as one of the solvers confirms the result to be optimal. Fig. 8 shows the cumulative distribution function over all used benchmark instances on selected problem characteristics. The vertical lines mark the average characteristic for instances where the best result was achieved by one solver (One Best), by a subset of the solvers (Multiple Best), and instances where all solvers could find the same best result (not necessarily the optimal result).
Table 2: Characteristics of all benchmarks.
mean std max min
$\beta$ 0.2806 0.1749 1.0 0.0667
$dv$ 0.1496 0.2241 0.95 5.6669e-05
N 150.1144 111.2573 500 4
$m$ 13.5124 13.1667 60 2
6.3.1 MILP
Gurobi
The characteristics of the instances Gurobi found the best solutions for are shown in Tab.3. Gurobi could not solve any instance that could not also be solved by at least one other solver. The instances that can be solved by Gurobi all reach the known optimum (11). The instances in question are small in terms of number of machines and number of operations. The flexibility is high in comparison to the average flexibility in all benchmarks (Tab. 2). The same is true for the duration variety.
Table 3: Characteristics of benchmarks in which Gurobi achieved the best result.
Average Benchmark Attributes
count $m$ N $\beta$ $dv$
mean std mean std mean std mean std
Single Best 0 - - - - - - - -
Shared Best 11 5.63 2.203 22.63 19.69 0.32 0.14 0.4486 0.3507
Combined 11 5.63 2.203 22.63 19.69 0.32 0.14 0.4486 0.3507
Optimal 11 5.63 2.203 22.63 19.69 0.32 0.14 0.4486 0.3507
CPLEX-LP
The characteristics of the instances the LP solver included in CPLEX found the best solutions for are shown in Tab. 4. The attributes are similar to the instances solved by Gurobi. While CPLEX-LP also could not solve any instances as the only solver, it managed to solve 11 instances to the known optimum. The instances solved by CPLEX-LP are smaller than the ones solved by Gurobi on average. The flexibility is slightly lower, while the duration variety is significantly higher.
Table 4: Characteristics of benchmarks in which the LP solver included in CPLEX achieved the best result.
Average Benchmark Attributes
count $m$ N $\beta$ $dv$
mean std mean std mean std mean std
Single Best 0 - - - - - - - -
Shared Best 11 4.18 1.83 15.73 14.21 0.32 0.21 0.6506 0.2539
Combined 11 4.18 1.83 15.73 14.21 0.32 0.21 0.6506 0.2539
Optimal 11 4.18 1.83 15.73 14.21 0.32 0.21 0.6506 0.2539
6.3.2 CP
OR-Tools
The characteristics of the instances OR-Tools found the best solutions for are shown in Tab. 5. OR-Tools managed to reach the best solution in 15 instances that could not be solved as well by any other solver. It also reached the best result for 56 instances for which at least one other solver reached the best result. Of the 71 best results reached, 48 are confirmed to be optimal. The average amount of machines is slightly higher than for both MLIP approaches, while the standard deviation is similar. Compared to CPLEX-CP and the GA the amount and spread of machines is low, but similar to Hexaly. For the amount of operations, OR-Tools solves bigger instances than both MILP approaches, but smaller than Hexaly, CPLEX-CP and the GA. The flexibility is smaller than the MILP approaches, CPLEX-CP, and GA, and similar to Hexaly. In comparison to Hexaly, CPLEX-CP and the GA, OR-Tools requires a higher duration variety to succeed, but can handle a lower duration variety than both MILP approaches, considering that the duration variety does not significantly change for the different categories shown in the table.
Table 5: Characteristics of benchmarks in which OR-Tools achieved the best result.
Average Benchmark Attributes
count $m$ N $\beta$ $dv$
mean std mean std mean std mean std
Single Best 15 7.47 2.45 78.0 29.33 0.11 0.06 0.3061 0.2607
Shared Best 56 6.11 2.96 53.02 40.21 0.39 0.25 0.3912 0.302
Combined 71 6.39 2.90 58.29 39.35 0.36 0.24 0.3732 0.294
Optimal 48 5.67 2.23 47.31 33.58 0.43 0.26 0.3960 0.296
CPLEX-CP
The characteristics of the instances CPLEX found the best solutions for are shown in Tab. 6. CPLEX-CP managed to find the best solution in 283 cases where no other solver found the best solution. The best result was also found for 66 additional benchmark instances for which other solvers also found the best solution. 69 of the overall 349 found best results are the known optimum of that benchmark instance. CPLEX-CP can handle lower duration varieties than the MILP approaches and OR-Tools, as well as slightly lower duration varieties than Hexaly. However, it can not handle the lowest duration varieties like the GA. In terms of the flexibility, the average flexibility of the instances only CPLEX-CP could solve best is slightly lower than the flexibility where the GA performs best, but higher than for all the other solvers. The average values for the instances that only CPLEX-CP solved best are very close to the average instance attributes (Tab. 2) of all benchmark instances. CPLEX-CP can handle a high operation amount in comparison to all other solvers except the GA. The average amount of machines is also higher than for the other solvers, the exception again being the GA.
Table 6: Characteristics of benchmarks in which CPLEX achieved the best result.
Average Benchmark Attributes
count $m$ N $\beta$ $dv$
mean std mean std mean std mean std
Single Best 283 14.14 12.55 175.96 112.49 0.28 0.15 0.1026 0.177
Shared Best 55 15.99 21.27 63.53 51.41 0.04 0.06 0.3424 0.294
Combined 338 14.20 13.89 158.14 114.98 0.014 0.03 0.1361 0.216
Optimal 69 20.77 21.83 73.99 62.06 0.039 0.05 0.3059 0.303
6.3.3 SBO
The characteristics of the instances the GA found the best solutions for are shown in Tab. 7. The GA could find the best solution for 31 instances, 17 of which are the known optimum. Any of the 7 best solutions that only the GA could find could also be the optimum, but the GA is lacking any method of confirming the optimum. Note that the stable value of 60 for amount of machines point at the Behnke instances [5], which contain the largest problem instances. For the instances that only the GA could solve best, the amount of machines and operations are high on average with a big range especially for the amount of operations. Paired with the flexibility values, which are slightly higher than CPLEX-CP and significantly higher than for the other solvers, this indicates that the GA is well suited for large and complex instances. Additionally, the GA performed better than all other solvers for instances with a low duration variety, which makes these instances harder to solve. Through the average metrics for the Shared Best instances, it can be seen that the GA can still also perform well on instances with high duration variety, and higher flexibility. Measuring the stability of the GA for every benchmark instance using Eq. (24), the average spread around the average result over all instances is 6.35% (median 5.01%).
$\begin{split}\delta=\cfrac{f_{worst}-f_{best}}{f_{average}}\end{split}$ (24)
For the instances for which the GA could find the best result overall, the average amount of machines stands out compared to all other solvers. The average amount of operations is also higher than the other solvers, though Hexaly comes close to a similar problem size, but with a much lower standard deviation. The average flexibility is below OR-Tools but above CPLEX-CP.
Table 7: Characteristics of benchmarks in which the GA achieved the best result.
Average Benchmark Attributes
count $m$ N $\beta$ $dv$
mean std mean std mean std mean std
Single Best 7 60.00 0.00 214.29 207.59 0.16 0.02 0.0003 0.0003
Shared Best 24 11.75 16.73 44.33 71.21 0.34 0.21 0.4393 0.358
Combined 31 22.65 25.20 82.71 133.13 0.29 0.19 0.3402 0.3648
Optimal 17 4.65 2.42 12.12 7.12 0.39 0.19 0.6181 0.2603
The characteristics of the instances for which Hexaly found the best solutions are shown in Tab. 8. Overall, Hexaly found the best solution for 71 of the instances, 37 of which were identified as the optimum (by any of the solvers). 29 of the best results were on instances where only Hexaly could find the best result, and 42 of the best results were also found by at least one other solver. Hexaly performs well on instances with a higher duration variety than the GA and CPLEX-CP but handles a lower $dv$ better than OR-Tools and the MILP approaches. Hexaly seems to deal best with the lowest flexibilities out of all solvers. The problem size is narrower spread than for the Single Best results of the GA and CPLEX-CP and on average lower, but higher than the other solvers. The same applies to the amount of machines.
Table 8: Characteristics of benchmarks in which Hexaly achieved the best result.
Average Benchmark Attributes
count $m$ N $\beta$ $dv$
mean std mean std mean std mean std
Single Best 29 10.10 1.92 116.76 52.25 0.07 0.05 0.1424 0.153
Shared Best 42 6.12 3.19 53.12 52.19 0.24 0.19 0.3887 0.2943
Combined 71 7.75 3.36 79.11 60.66 0.17 0.17 0.2881 0.2738
Optimal 37 5.43 2.15 50.27 54.41 0.26 0.19 0.3984 0.289
6.4 Comparison
Figure 9: The portion of problem instances per solver that are within a relative gap (Eq. (23)) to the best known result.
Optimal Feasible Infeasible
CPLEX_CP 69 331 2
CLPLEX_LP 11 310 81
Gurobi 11 305 86
Hexaly 37 365 0
OR-Tools 48 349 5
GA_Best 17 385 0
GA_Average 14 388 0
Table 9: Caption
The results in Fig. 9 show the relative gap of the results of each solver to the best result found between all solvers. While the results in [12] showed that MILP solvers perform better on the FJSSP instances, the results indicated that the GA would start to outperform the MILP solver on more complex problems. This can be shown through the data collected using the FJSSP-W instances. Considering the best results of the GA as well as the average results of the GA perform better than both used MILP solvers for the FJSSP-W. This shows that the GA can scale to more complex scenarios better. Fig. 9 shows that except for the MILP approaches, all solvers are less than 25% worse than the best overall found result on over 80% of all instances.
Given even more complex scenarios, like the FJSSP-W with added uncertainties for robust optimization, the GA can be expected to increase it’s advantage over the other solvers even further.
As seen in the results (see Sec. 6.3), all solvers show tendencies towards a certain category of problem instance. While the GA seems to scale best to large, complex problems, the CP solvers are closest to solving the average benchmark instance for the given benchmark collection. The known optima found by the solvers all have a higher than average duration variety, showing that a high duration variety indicates instances which are easier to solve across all solvers. The same seems to be true for the flexibility. Considering the combination of high duration variety and high flexibility, this means that for those instances the assignment problem can be solved easier, leaving the main pressure on solving the sequencing problem. However, the average amount of machines and operations on instances that are solved optimally is lower than the average benchmark instance, meaning that mostly small instances with high duration variety are the ones that can be solved by any of the solvers easily.
Table 10: Average characteristics of optimally solved benchmarks.
mean std
$\beta$ 0.4145 0.2251
$dv$ 0.03872 0.0542
N 73.9855 62.0568
$m$ 7.1739 3.3429
$d_{mean}$ 100 131.3469
$d_{stdv}$ 52.9665 79.3247
$d_{span}$ 189 272.8644
Figure 10: Benchmark attributes of all instances normalized over the maximum value.
The characteristics of the benchmarks that could be solved optimally are shown in Tab. 10. From this data it can also be seen that the solvers generally seem to profit from an easier assignment problem, with a low number of machines on average and a mid-range flexibility. For these smaller instances, a smaller duration variety also appears to be less of an issue, for the same reason. The distribution of actual duration values was also considered (Fig. 10) , since some problem instances deal with duration values smaller than 10 while other instances have duration values up to over 1000, with he maximum span present in the benchmark set being over 1600. However, due to the large variation of duration value metrics in the optimally solved instances (Tab. 10), the actual values do not seem to impact the solvers, as long as the duration variety decreases the number of solutions with the same fitness values and the amount of mutations applied to solutions that do not lead to change in their fitness value. In comparison to the average characteristics of all benchmarks (Tab. 2), the flexibility of the optimally solved instances is higher and the problem size is lower.
Overall the data implies that the assignment part of the optimization (machines for FJSSP, machines and workers for FJSSP-W) seems more problematic for all solvers but the GA, which additionally explains why the GA scales up better to the FJSSP-W than other solvers.
For further investigation, it could prove useful to craft a better balanced benchmark collection, varying in problem size (machines, workers, operations) as well as their complexity (flexibility, duration variety), to cover the metrics more evenly for testing and comparison purposes. Many of the different benchmark sources also contain multiple instances that cover very similar problems, which results in a lot of experiment runtime is used to test the solvers on the essentially the same problem.
(a)Refer to caption (b)Refer to caption (c)Refer to caption
Figure 11: Best solver for benchmark characteristics
(a)Refer to caption (b)Refer to caption (c)Refer to caption
Figure 12: Optimal solution found by the different solvers
(a)Refer to caption (b)Refer to caption (c)Refer to caption
Figure 13: Best solution based on solver category. While MILP approaches could not achieve the single best result for any instance, they could optimize 11 instances the same as all other solvers. CP approaches achieved by far the most singled out best results (315), followed by Hexaly (29), then the GA (8). 39 instances were solved best by a mixture of solver categories and 11 instances were solved the same by all solver categories.
7 Conclusion
While the GA only outperforms the MILP approaches on the FJSSP on specific instances  [12], the MILP approaches are not capable of scaling up to the FJSSP-W as well as the GA. Except for CPLEX-CP, the same is true for the other approaches, like the CP approach using OR-Tools and Hexaly. While both can still keep up with the GA on certain instances, the GA shows better potential in scaling up to more complex scenarios. In direct comparison between the CP approach and the GA approach, the characteristics of the benchmark problems solved best by either show some significant differences. Namely, the GA deals better with a more difficult assignment problem, with both more machine and workers as well as more operations. However, for simpler assignment problems (low $dv$), the CP approach outperforms the GA in runtime. While it is likely that the efficiency of the implementation plays a role for those problems, the CP approach does tend to find its final solution earlier on in the optimization process than the GA, which steadily increases its best solution over time. This shows that both approaches have an advantage in different scenarios, which can possibly be determined by the problem characteristics, and could be exploited by a hybrid model. The capability of the GA to handle low $dv$ values in large problem instances demonstrates its suitability for even more complex use-cases, including FJSSP and FJSSP-W with uncertainties.
8 Outlook
To make the current system more applicable to the real-world, the developed solvers need to be more robust towards uncertainties. Given that the GA shows stability when scaled to more complex problems, it is well geared to include uncertainty in its optimization process, while other approaches like MILP and CP struggle with uncertainty. However, given the efficiency of CP to find good solutions, a hybrid method combining the advantages of both CP and GAs will be considered as well. Different types of uncertainties will be included in the future work, like unexpected events (e.g. machine breakdowns, worker unavailability) and inaccurate information (e.g. stochastic processing times).

### v2

@@
-Comprehensive Benchmarking Environment for Worker Flexibility in Flexible Job Shop Scheduling Problems
+A Benchmarking Environment for Worker Flexibility in Flexible Job Shop Scheduling Problems
 Abstract
@@
 Hence, each machine also requires a worker to be present to process an operation which additionally affects the processing times.
-The resulting problem is called Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W).
-The FJSSP has been approached with different strategies considering various problem representations, including Mixed Integer Linear Programming (MILP), Constrained Programming (CP), and Simulation-based Optimization (SBO).
-In the latter area in particular, there exists a large number of specialized Evolutionary Algorithms (EA) like Particle Swarm Optimization (PSO) or Genetic Algorithms (GA).
-Yet, the solvers are often developed for single use cases only, and validated on a few selected test instances, let alone compared with results from state-of-the-art solvers using other problem representations.
-While suitable approaches do also exist, the design of the FJSSP-W instances is not standardized and the performance of the algorithms is hardly comparable.
-This calls for a systematic benchmarking environment that provides a comprehensive set of FJSSP(-W) instances and supports targeted algorithm development.
-It will facilitate the comparison of algorithmic performance in the face of different problem characteristics.
+The resulting problem is called Flexible Job Shop Scheduling Problem with Worker Flexibility.
+Such problems are approached with different strategies considering various problem representations, including Mixed Integer Linear Programming, Constrained Programming, and Simulation-based Optimization.
+In the latter area, in particular, there exists a large number of specialized Evolutionary Algorithms.
+Yet, solvers are often developed for distinct use cases only, and validated on few selected test instances, let alone compared with results from multiple state-of-the-art solvers.
+While suitable approaches exist, the design of the FJSSP-W instances is not standardized and algorithm performance is hardly comparable.
+This requires a systematic benchmarking environment that provides a comprehensive set of FJSSP(-W) instances for targeted algorithm development.
+It will facilitate comparing algorithmic performance in the face of different problem characteristics.
 The present paper presents a collection of 402 commonly accepted FJSSP instances and proposes an approach to extend these with worker flexibility.
-The conversion of the FJSSP into an FJSSP-W has a significant impact on problem complexity.
-In addition, we present a detailed procedure for the evaluation of scheduling algorithms on these problem sets and provide suitable model representations for this purpose.
-This includes MILP formulations, CP models, and appropriate problem encodings for SBO solvers for both problem types.
-In addition, we provide complexity characteristics for all presented instances as well as baseline results of common commercial solvers to facilitate the validation of new algorithmic developments.
+Additionally, it presents an environment for the evaluation of scheduling algorithms on these problem sets.
+Further, instance characteristics and baseline results of common commercial solvers are presented to facilitate the validation of new algorithmic developments.
 keywords:
@@
 […]
-[inst1]organization=Josef Ressel Centre for Robust Decision Making
-Research Centre Business Informatics
-Vorarlberg University of Applied Sciences,addressline=
-Hochschulstrasse 1, city=Dornbirn, postcode=6850, country=Austria
+[inst1]organization=Research Centre Business Informatics, Josef Ressel Centre for Robust Decision Making,
+Vorarlberg University of Applied Sciences,addressline=Hochschulstrasse 1, city=Dornbirn, postcode=6850, country=Austria
 {highlights}
-An alternative modeling approach for flexible job shop scheduling problems (FJSSP) as a mixed integer linear program (MILP) suitable to represent an exhaustive set of commonly accepted benchmark problems is proposed.
-The paper presents a refined algorithmic approach to extend the FJSSP model with worker flexibility constraints (FJSSP-W).
-This way, we provide a portfolio of benchmark instances for an otherwise rather fragmented problem branch.
-A corresponding extension of the derived MILP representation for the introduced FJSSP-W is also delivered.
-We provide a suitable problem encoding for a convenient simulation-based optimization (SBO) via meta-heuristics, e.g.
-Evolutionary Algorithm (EA) variants, to such problem classes and present an approach to effectively treat the worker flexibility constraints.
-In addition, we develop adequate Constraint Programming (CP) representations for both problem classes (FJSSP and FJSSP-W) which facilitate the comparison with state-of-the-art solvers from this field of research.
-Two problem classes in three representations (MILP, CP, and SBO) provide a comprehensive test environment for the comparison of solvers of different algorithmic paradigms.
-As a baseline, we compare the results of commercial state-of-the-art solvers (GUROBI, Hexaly, CPLEX (IBM), OR-Tools (Google)) and self-developed Genetic Algorithms (GA).
+Introduces an algorithmic method to enhance FJSSP with worker flexibility constraints.
+Establishes a portfolio of benchmark instances for FJSSP with Worker Flexibility (FJSSP-W).
+Presents an open-source benchmarking environment for FJSSP and FJSSP-W solvers.
+Supports algorithm development with user-specific filtering and harmonized results.
+Provides problem encoding for multiple optimization paradigms: meta-heuristics, MILP, and CP.
 1 Introduction
-Planning problems occur in a variety of application areas where the allocation and optimization of resources is critical to achieving specific goals.
-Consequently, planning problems are relevant in virtually all economic areas and the corresponding challenges are critical for ensuring sustainable economic growth, competitiveness, and efficient operations.
-In Supply Chain Management, for example, planning is essential for coordinating the flow of goods.
-In the Energy Sector, planning problems deal with the generation, distribution, and consumption of energy.
-Other application fields include Healthcare, Finance, Agriculture, and many more.
-Another important area is Manufacturing, i.e.
-the branch of industry that involves the production of finished products from raw materials.
-There, Production Scheduling (PS) deals with the allocation of resources such as machinery, labor and materials in order to achieve production targets efficiently.
-Such production goals include meeting customer requirements in terms of delivery times, minimizing idle time in certain production steps or responding to machine breakdowns.
-In this way, effective scheduling aims to promote sustainable development, drive innovation and ensure the resilience and competitiveness of manufacturing companies.
+Planning problems arise in various fields where resource allocation and optimization are crucial for achieving goals.
+These problems are relevant across all economic sectors, impacting sustainable growth, competitiveness, and efficiency [1, 2, 3].
+In Supply Chain Management, planning coordinates the flow of goods.
+In the Energy Sector, it addresses the generation, distribution, and consumption of energy.
+Other fields include Healthcare, Finance, and Agriculture.
+In Manufacturing, Production Scheduling (PS) allocates resources like machinery, labor, and materials to meet production targets efficiently.
+Goals include meeting delivery times, minimizing idle time, and responding to machine breakdowns.
+Effective scheduling promotes sustainable development, drives innovation, and ensures the resilience and competitiveness of manufacturing companies.
 Production Scheduling problems involve determining the proper sequence of actions or decisions to attain (near) optimal performance on the shop floor.
-They take a variety of forms and are commonly approached by applying techniques from Artificial Intelligence (AI) or Operations Research (OR).
+They take a variety of forms and are commonly approached by applying techniques from Artificial Intelligence (AI), Operations Research (OR) or Simulation-based Optimization (SBO) [2].
 A broad overview of the PS landscape is provided in Fig.
 1, in which the problem branches on which this work is based are indicated by the blue boxes.
-Figure 1: Localization of the problem classes under consideration in the area of Production Scheduling.
+Figure 1: Visualization of the problem classes in the area of Production Scheduling considered in this paper.
 Notice that other problem types exist and the illustration does by no means claim to be comprehensive.
-Among such tasks, the Job Shop Scheduling Problem (JSSP) is a type of scheduling problem where multiple jobs (or tasks) share a common set of resources but have different processing requirements and routes through the production facility.
-Each job may require different machines or processes, and the order in which the jobs are processed can affect the overall efficiency of the system.
-JSSPs may turn out very complex due to the variability in job processing times, resource constraints, and the need to optimize the sequence of operations.
-Unlike the traditional JSSP, Flexible Job Shop Scheduling Problem (FJSSP) allows jobs to be processed on different machines.
-This provides more flexibility in resource allocation to those machines that possesses the required capabilities for a job.
-In this respect, the FJSSP represents an extension of the JSSP that considers a more dynamic and adaptable environment.
-The flexibility allows for more dynamic and adaptive scheduling, but it also introduces additional complexity to the scheduling problem, and optimization algorithms need to consider the various combinations of machine assignments to achieve an (near) optimal schedule.
-An additional layer of complexity to the scheduling process is added when worker-related dynamics are also considered within a FJSSP.
-Such problems are commonly referred to as Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W).
-There are many restrictions that may arise when considering worker flexibility in FJSSP, e.g., consecutive working hours, breaks or different qualifications of workers.
-Depending on the modeling approach used and the optimization technique envisaged, JSSP and FJSSP model representations come in different forms.
-Common problem formulations include mathematical graphs, such as Resource-Task-Network (RTN) representations, permutation-based representation, or mathematical programming (MP) formulations in the form of objective functions and constraints.
-The latter usually come in the form of Mixed-Integer Linear Programming (MILP) problems and can be treated by various optimization techniques.
-Formulations as simulation-based optimization (SBO) problems are also frequently found in the literature.
-Instead of an analytical representation of the problem, the objectives and dependencies of a production line are implemented in the form of a black-box environment that allows the evaluation of individual production plans.
-Constraint Programming (CP) formulations are another way of expressing production planning problems.
-CP enables the integration of constraints and their solution mechanisms into a programming language.
-The problem is formulated in declarative form, while the solution process is secondary from the users’ point of view.
-This task is performed by a specialized constraint solver.
-For SBO representations common approaches comprise meta-heuristic algorithms that rely on nature-inspired working principles, such as Genetic Algorithm (GA), Ant Colony Optimization (ACO) or Particle Swarm Optimization (PSO) variants.
-Hybrid approaches that combine multiple optimization techniques to exploit complementary strengths are also regularly considered.
-An example would be the combination of a GA with a Local Search method.
-Focusing on the exploration of the neighborhood of a given solution, Local Search (LS) solvers are used to refine solutions obtained from other methods and to find locally optimal solutions.
-In case of a MILP or CP formulation of the FJSSP, specialized commercial mathematical programming solvers of e.g.
-GUROBI and CPLEX (IBM) or constraint programming solvers like OR-Tools(Google), Hexaly or CP of CPLEX respectively, are widely used to find (near) optimal solutions.
-These solvers obtain state-of-the-art solutions for many FJSSP instances.
-Still, the choice of solver depends on factors such as the size and complexity of the FJSSP instance, the desired solution quality, and the available computational resources.
-Researchers and practitioners often experiment with multiple approaches to find the most effective solution for a specific FJSSP instance.
-While the performance of most solvers is assessed on individual real-world problem instances and compared to selected solver variants only, some scheduling approaches are validated on rather small sets of test problem instances.
-In contrast to the real-world problems considered, the majority of FJSSP test sets must be considered fairly academic.
-Nevertheless, recent literature offers access to a number of standard FJSSP benchmark instances being spread across individual publications and modeled in different ways.
-Further, the benchmark problems usually vary in their structure and complexity.
-A unified benchmarking framework which allows dedicated algorithm performance comparisons on FJSSPs does not exist to the best of our knowledge.
-Moreover, there are hardly any comparable benchmark instances for FJSSP-W.
-As the modeling of FJSSP-W requires a careful consideration of both machine and worker-related constraints.
-The type of model affects the selection of solvers able to handle the increased complexity introduced by flexible scheduling policies.
-This introduces novel objectives like not only optimizing production efficiency but also taking into account the necessities of the workforce.
-This paper proposes a systematic test environment of FJSSP and FJSSP-W instances to benchmark state-of-the-art solvers.
-For this purpose, an alternative MILP modeling approach for FJSSPs suitable for incorporation of a comprehensive set of 402 commonly accepted benchmark problems is introduced.
-Our MILP formulation allows for an improvement of the schedules found by the commercial solver GUROBI on some of the benchmark instances.
-The paper presents an algorithmic approach to extend the FJSSP model with worker flexibility constraints (FJSSP-W).
-This way, we provide a way to systematize the fragmented area of FJSSP-W instances.
-It provides an opportunity to create a portfolio of benchmark instances for the rather fragmented area of FJSSP-W instances.
-Further, suitable SBO problem encodings convenient for the application of meta-heuristics, e.g.
-Evolutionary Algorithm (EA) variants, as well as adequate CP representations are provided.
-As a result, a comprehensive test environment for the comparison of solvers from three different paradigms (MILP, CP, and SBO) on both problem classes (of 402 instances each) is established.
-Lastly, the paper describes a structured benchmark environment for the FJSSP(-W) which offers the possibility to choose among different problem characteristics, as well as the evaluation of a set of performance indicators (relative solution quality, wall clock time, etc.) and corresponding graphical visualizations.
-The test environment also provides characteristic attributes for each problem instance, such as the total number of operations, the flexibility of the machines or the problem-related duration variety, on the basis of which a targeted selection of sub-instances can be made.
-The functionalities are demonstrated by presenting baseline results of various state-of-the-art commercial solvers (GUROBI, Local Solver, CPLEX (IBM), OR-Tools (Google)) and a self-developed Genetic Algorithms (GA).
-The remainder of this paper is structured as follows: In Section 2, a detailed description of the considered problem classes is provided.
-In addition, related work is discussed and available benchmarks as well as state-of-the-art solvers are presented.
-Section 3 deals with the formulation of the respective FJSSP representations.
-Next, the procedure for integration of worker flexibility is specified in Sec.
-4.
-It realizes the transformation of an FJSSP instance into an FJSSP-W instance with just a few user-defined assumptions.
-Subsequently, the corresponding FJSSP-W formulations are specified.
-The detailed description of the benchmarking environment and its configuration options are the subject of Sec.
-5.
-Afterwards, Section 6 presents a number of baseline results obtained by application of several state-of-the-art solvers.
-The contribution of this work and its impact on the field of research is discussed in summary in Sec.
-7.
-Finally, Section 8 gives an outlook on future extensions of the presented benchmarking suite.
+The Job Shop Scheduling Problem (JSSP) involves scheduling multiple jobs that share resources but have different processing requirements and routes.
+Each job may need different machines or processes, and the order of processing affects system efficiency.
+JSSPs are complex due to variable job processing times, resource constraints, and the need to optimize operation sequences.
+Unlike the classical JSSP, Flexible Job Shop Scheduling Problem (FJSSP) allow jobs to be processed on different machines.
+This provides more flexibility in resource allocation and may be regarded an extension of the JSSP that considers a more dynamic and adaptable environment.
+Yet, the flexibility introduces additional complexity, as optimization algorithms need to deal with the combinations of machine assignments to achieve an (near) optimal schedule.
+In case that workforce-related dependencies are included, the problem is known as Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W).
+This adds complexity due to factors like consecutive working hours, breaks, and worker qualifications on the individual machines.
+A comprehensive collection and descriptions of commonly used FJSSP instances are provided in [4].
+The paper also proposes some additional FJSSP benchmark instances and provides best-known results for all reported instances.
+In 2024, a more recent survey [3] discusses common problem instances, different problem constraints, as well as solution approaches.
+While the paper updates the best-known results for the discussed FJSSP instances with more recent findings, it must be noted that there are problem instances in both works that are not considered in the respective other work.
+The paper concludes that the most important optimization criterion is the time span [3], although other criteria are also discussed.
+These include machine utilization, delays, weighted lead times, energy costs, queue times and just-in-time measures.
+Additional constraints such as minimum waiting times between two operations, a maximum waiting time for starting orders in the schedule, resource availability constraints, batch formation constraints, setup time constraints, resource blocking constraints (e.g.
+limited queue space at the machines), transportation constraints and some other special cases can be found in the literature, but they do not occur to a significant extent.
+As far as the systematic benchmarking [5, 6] of algorithmic solvers for such PS problems is concerned, only a few partial aspects are covered, particularly in the area of SBO methods.
+The development and the performance comparison of of such solvers particularly relies on benchmarking due to the lack of theoretical performance results for optimization tasks of notable complexity.
+First and foremost, benchmarking experiments are established for performance evaluation and algorithm comparison on given problem classes.
+Ideally, this is supposed to support the selection of the algorithm best suitable for given real-world applications [7].
+Further, benchmarks can be used to deliver insights into the working principles of algorithms and foster their development for specific problem types [8, 9].
+Unlike other problem areas such as continuous single objective constrained optimization problems [10] or combinatorial Traveling Sales Person problems [11], to the best of our knowledge, there is no commonly accepted benchmarking environment for the FJSSP (Flexible Job Shop Scheduling Problem) for the aforementioned purposes.
+Instead, the situation is segmented into the following categories:
+Sets of problem instances
+Over the years, multiple authors proposed their own versions of FJSSP test problem instance sets [4, 12].
+These problem sets are widely accepted and used for the benchmarking of algorithmic developments.
+In addition, new problem instances are constantly being created [13].
+Although, comprehensive collections of such sets are openly available [14, 15, 16, 4], the rather arbitrary use of instance subsets is the usual practice.
+This complicates the cross-comparability of solvers and impedes a preferably complete picture of the situation.
+Real-world problem instances
+There are many individual real-world problems that are used to justify algorithmic developments.
+While there are efforts to derive reasonable benchmarking environments form specific industrial use cases [17, 18], the majority of cases does not allow for the (systematically) comparison of different solvers.
+This is partly due to the real-world problems coming with very individual requirements that complicate the application of many solvers [19].
+On the other hand, the computational effort can be too extensive for a rigorous assessment.
+Moreover, not all details of the real-world problem dependencies might be properly disclosed in all cases [18].
+Competitions
+In the context of conferences such as GECCO [^1] An example is the GECCO2025 Competition on Evolutionary Computation in the Energy Domain: Summer Finals of the Risk-based Energy Scheduling which is held during the Genetic and Evolutionary Computation Conference in 2025 https://www.gecad.isep.ipp.pt/ERM-competitions/2025-2/.
+and CEC [^2] For example, in the context of the Congress on Evolutionary Computation, the CEC’2024 Competition on Large-Scale Sequence Optimization for Production Lines was organized http://developer.xuelangyun.com/#/competition?competitionId=1009000&lang=en.
+In 2025, the same conference hosts the CEC’2025 Competition on Dynamic Scheduling of Flexible Flow Shop with Random Job Arrivalshttps://github.com/SIA-aps/CEC2025-DFFSP-competition.
+, competitions on scheduling problems are regularly announced.
+These are usually well documented and structured so that they can also be used for benchmarking purposes.
+However, these competitions are usually also tailored to specific use cases from the areas of production or planning, sometimes with very special requirements.
+A broader comparison with solvers that have not been adapted to these problem-specific particularities is therefore difficult.
+Performance metrics
+A frequently discussed benchmarking aspect in Scheduling Problems is the definition and selection of performance metrics [20, 21].
+These can vary greatly depending on the scheduling objective, application area and other influencing factors (noise, dynamic constraints, etc.) [22].
+While it is hardly possible to prescribe a uniform guideline in this respect, it is important to pay close attention to the consistent use of such performance measures when comparing algorithms.
+When it comes to considering worker-related constraints, i.e.
+the FJSSP-W, the benchmarking landscape becomes even more scattered.
+There is neither a commonly accepted set of FJSSP-W instances nor a standardized method for creating FJSSP-W test instances.
+Instead, researchers often develop test problem instances independently [23, 24] or rely on solitary problem representations from the real world [25, 26].
+The lack of standardization complicates benchmarking principles such as comparability and reproducibility, as documentation may be incomplete or not all aspects of the real-world problem are disclosed.
+Consequently, this paper presents a benchmark environment that summarizes a comprehensive collection of problem instances for both the FJSSP as well as the FJSSP-W and allows for a systematic assessment of corresponding solvers.
+To this end, those FJSSP instances from the collections of [4] and [3] are selected that fulfill the FJSSP definition and must not be assigned to more specialized problem classes (i.e.
+parallel machines, or flow store scheduling problems, cf.
+Fig.
+1).
+Such FJSSP instances are then extended to FJSSP-W instances.
+Through the traceable modification, the increase in complexity between the problem classes and its effects on algorithm performance can be better understood.
+The benchmark environment represents an open-access platform for the performance evaluation and comparison of scheduling algorithms for the problem classes mentioned.
+By specifying standardized test criteria and termination conditions, benchmarking can be guaranteed to be as harmonized as possible.
+This is accompanied by a coherent visualization of the benchmarking results.
+The platform will affect the relative performance assessments of state-of-the-art solvers and be useful for the selection of solvers for specific problem classes.
+Additionally, it supports the development of algorithms for specific problem characteristics, such as particularly high flexibilities in the production environment, through optional user-specific filtering based on the characteristics of the provided benchmark instances.
+By selecting a smaller number of representative instances for a particular problem property, algorithmic developments can be tested repeatedly with less computational effort before they have to be applied to the entirety of the instances.
+The benchmarking environment is designed primarily for the comparison and development of simulation-based optimization approaches, e.g.
+Evolutionary Algorithms (EA), and is nevertheless intended to be accessible for classical Mathematical Programming methods.
+To this end, the supplementary material of this paper includes Mixed-Integer Linear Programming (MILP) and Constraint Programming (CP) problem representations.
+These models also allow the reproducibility of the best-known results available in the benchmarking environment, which were obtained by applying multiple state-of-the-art solvers like Gurobi, Hexaly, CPLEX, or OR-Tools [27, 28, 29, 30].
+The paper is organized as follows: Section 2 elaborates on the FJSSP formulation and provides an overview of available benchmark sets.
+The extension towards FJSSP-W is then specified in Sec.3.
+In Sec.4, the benchmarking environment is introduced.
+This is followed by the demonstration of sample results of several test solvers in Sec.5.
+Finally, the paper concludes in Sec.6, and proposes some directions for future work in Sec.7, respectively.
 2 Problem Formulation & Related Work
@@
 In this section, a formal definition of this problem is presented in Sec.
-2.1 and current work on the FJSSP is outlined in Sec.
-2.3.
-A large proportion of such research looks at very individual problems from real-world applications or only compares the developed solvers on individually selected benchmark instances.
-This complicates the performance comparison between the solvers significantly.
-In order to counteract this situation and to develop a structured benchmarking environment, Sec.
-2.2 provides an overview of the available benchmark instances in the field of FJSSP and their characteristics.
+2.1 together with the introduction of necessary problem features (i.e.
+flexibility or duration variety) that allow for the characterization of individual problem instances.
+Further, Sec.
+2.2 provides an overview of the available benchmark instances in the field of FJSSP.
 2.1 Problem Formulation
@@
 The machines are not necessarily homogeneous which allows for different processing times for each operation, depending on which machine it is assigned to.
-In general formulation, the FJSSP introduces a set of $n$ jobs $\mathcal{J}=\{J_{1},\dots,J_{n}\}$ that need to be processed on a set of $m$ machines =$\mathcal{M}=\{M_{1},\dots,M_{m}\}$.
+Using the notation from [32], the general representation of a FJSSP introduces a set of $n$ jobs $\mathcal{J}=\{J_{1},\dots,J_{n}\}$ that need to be processed on a set of $m$ machines =$\mathcal{M}=\{M_{1},\dots,M_{m}\}$.
 Each of these jobs $J_{i}\>(i=1,\dots,n)$ consists of a predetermined sequence of $n_{i}$ operations.
 To this end, the notation $O_{i,j}$ refers to the $j$th operation within job $i$.
-Each of these operations $O_{i,j}$ is performed by one of the admissible machines in the set $\mathcal{M}_{i,j}\subset\mathcal{M}$.
+Each of these operations $O_{i,j}$ is performed by one of the $m_{i,j}$ admissible machines in the set $\mathcal{M}_{i,j}\subset\mathcal{M}$.
 If operation $O_{i,j}$ is assigned to a specific machine, this machine is denoted by $M_{i,j}$.
@@
 In common scenarios, the general representation of an FJSSP assumes that all orders are known and all machines are operational at time 00.
-In the context of FJSSPs, a large number of objective functions are conceivable with respect to the optimization of which the appropriate schedule must be composed.
+In the context of FJSSPs, many objective functions are conceivable, the optimization of which determines the final production plan.
 The performance criteria include minimizing the makespan, minimizing tardiness or maximizing resource utilization, to name but a few.
@@
 Note, that concrete examples of a parameter vector encoding for $\mathbf{y}$ is depending on the modeling approach and will be provided below.
-As it is common in the field of meta-heuristic optimization, the expression fitness will be used synonymously for the quality of a discovered feasible solution $\mathbf{y}$, i.e.
+As it is common in the field of meta-heuristic optimization, the term fitness will be used synonymously for the quality of a discovered feasible solution $\mathbf{y}$, i.e.
 a feasible schedule, in the subsequent sections of this work.
-That is, solution candidates of higher fitness (or quality) correspond to solutions of lower cost function values.
+Hence, candidate solutions of higher fitness (or quality) correspond to solutions of lower cost function values.
 The quality of two admissible scheduling solutions can consequently be compared on the basis of their objective function values.
 Accordingly, lower cost function values $C(\mathbf{y})$ correspond to better production plans $\mathbf{y}\in\mathcal{Y}$.
-For a proper comparison of two specific schedules, an order relation $\prec_{c}$ is defined by
+For a the comparison of two feasible schedules $\mathbf{y}$ and $\mathbf{z}$, an order relation $\prec_{c}$ is defined by
 $\mathbf{y},\mathbf{z}\in\mathcal{Y}:\qquad\mathbf{y}\prec_{c}\mathbf{z}\>\ratio\Leftrightarrow\>C(\mathbf{y})<C(\mathbf{z}).$ (1)
-Hence, the fitness of $\mathbf{y}$ would be rated higher than that of $\mathbf{z}$ in the case of Eq. (1). The corresponding smaller or equal relation $\preceq_{c}$ is defined analogously.
-2.2 Existing FJSSP Benchmarks
+That is, in this case the fitness of $\mathbf{y}$ is rated higher than that of $\mathbf{z}$ with respect to the cost function $C$ (e.g.
+the makespan of the schedules).
+The corresponding smaller or equal relation $\preceq_{c}$ is defined analogously.
+2.2 Foundational FJSSP benchmark problem instances
 This section discusses the existing benchmark instances for the Flexible Job Shop Scheduling Problem.
@@
 A majority of solvers are rather evaluated on problem subgroups originating from distinct sources, which means that the results of such studies are always subject to a latent problem selection bias.
-A large number of papers also only consider individual real-world problem representations.
+A large number of papers also only consider individual real-world problem representations [2].
 The algorithms developed in this way are highly specialized and are rarely compared carefully with established solvers.
@@
 for instances of relatively low complexity, the computing time required to achieve the optimum would be of greater interest.
-On the other hand, improvements of the solution quality are desirable as long as the exact optimum is unknown.
+On the other hand, improvements of the solution quality are desirable if the exact optimum is unknown.
 In this work, all $402$ distinct FJSSP instances of the ten test problem collections in Table 1 are considered.
-The table also indicates the number of instances (# instances) contained in each collections as well as the average number of operations per job (# avg.
-ops per job).
-The latter is within the range of $3$ and $19.5$, indicating the differences between the proposed benchmarking sources.
+The table also indicates the number of instances (# inst) contained in each collections as well as other average characteristics, e.g.
+the average number of jobs per instance $\bar{n}$, or the average number of machines $\bar{m}$.
+The average number of operations per job $\bar{\texttt{N}}/\bar{n}$ lies within the range of $2.95$ and $19.49$, indicating the differences between the proposed benchmarking sources.
 Apart from the benchmark sources in Table 1, other collections with close relationships must be named.
@@
 Table 1: An overview of ten FJSSP benchmark problem collections from literature.
-No. Source name # instances # avg. ops per job Reference
-1 BehnkeGeiger 60 5,0 [5]
-2 Brandimarte 15 8,6 [6]
-3 ChambersBarnes 21 11,7 [7]
-4 DPpaulli 18 19,5 [8]
-5 Fattahi 20 3,0 [9]
-6 Kacem 4 3,2 [10]
-7 Hurink Edata 66 8,9 [11]
-8 Hurink Sdata 66 8,9 [11]
-9 Hurink Rdata 66 8,9 [11]
-10 Hurink Vdata 66 8,9 [11]
+No. Source name # inst $\bar{n}$ $\bar{\texttt{N}}$ $\nicefrac{{\bar{\texttt{N}}}}{{\bar{n}}}$ $\bar{m}$ $\bar{\beta}$ $\bar{dv}$ Ref.
+1 BehnkeGeiger 60 45.00 225.00 5.0 40.0 0.316 0.009 [4]
+2 Brandimarte 15 20.33 171.87 8.56 9.13 0.310 0.007 [35]
+3 ChambersBarnes 21 13.33 158.33 11.67 13.667 0.089 0.007 [36]
+4 DPpaulli 18 15.00 292.00 19.49 7.667 0.330 0.004 [37]
+5 Fattahi 20 5.35 17.40 2.95 5.1 0.517 0.098 [38]
+6 Kacem 4 9.75 31.75 3.16 8.0 1.0 0.042 [39]
+7 Hurink EData 66 14.76 133.38 8.85 8.848 0.151 0.010 [40]
+8 Hurink SData 66 14.76 133.38 8.85 8.848 0.131 0.010 [40]
+9 Hurink RData 66 14.76 133.38 8.85 8.848 0.258 0.010 [40]
+10 Hurink VData 66 14.76 133.38 8.85 8.848 0.476 0.010 [40]
+Overall 402 18.89 150.11 8.54 13.510 0.281 0.014
 The usual specification of the production environments is carried out for each individual FJSSP instance in the form of a text file, which shows all the necessary information about the dependencies in the flexible job shop.
@@
 The first row of such a file commonly contains two (optionally three) numbers: the first value determines the number of jobs $n$ and the second one specifies the number of machines $m$ in the FJSSP instance.
-An optional 3rd value may indicate the average number of machines per operation.
+An optional third value may indicate the average number of machines per operation.
 Although the latter is not necessary for the implementation of the problem instance.
@@
 The first value in each row gives the number of operations $n_{i}$ to be executed in job $J_{i}$.
-The second number is the number of machines $\lvert\mathcal{M}_{i,1}\rvert$ available to process the first operation $O_{i,1}$.
-This is followed by a sequence of $\lvert\mathcal{M}_{i,j}\rvert$ pairs of numbers $(k,T_{i,1,k})$ indicating the index $k$ of the admissible machines $M_{k}$ as well as the processing times $T_{i,1,k}$ of operation $O_{i,1}$ on machine $M_{k}$.
-Next, the data for the second operation $O_{i2}$ are provided.
+The second number is the number of machines $m_{i,j}$ available to process the first operation $O_{i,1}$.
+This is followed by a sequence of $m_{i,j}$ pairs of numbers $(k,T_{i,1,k})$ indicating the index $k$ of the admissible machines $M_{k}$ as well as the processing times $T_{i,1,k}$ of operation $O_{i,1}$ on machine $M_{k}$.
+Next, the data for the second operation $O_{i,2}$ are provided.
 In this way, all operations of all jobs are specified accordingly.
-In order to facilitate a systematic classification of the 402 FJSSP test instances and consequently support an assessment of the evaluation results, the characteristic properties for each instance are given in the Appendix.
+To facilitate a systematic classification of the 402 FJSSP test instances and consequently support an assessment of the evaluation results, the characteristic properties for each instance from each source are given in a public repository on Github [41].
 In addition to the key properties such as the total number of jobs ($n$), machines ($m$), or operations (N) of an instance, seven other instance-specific statistics are provided that allow for a rough intuition of the individual benchmark instance complexity.
 These include the minimum ($T_{\min}$), maximum ($T_{\max}$) and average processing times ($\bar{T}$) measured over all possible machine-operation combinations, as well as the corresponding standard deviation $\sigma_{T}$.
-Moreover, the average number of operations per job $\frac{\texttt{N}}{n}$, the machine flexibility $\beta$, and the duration variety $dv$ are provided.
-The latter two indicators are used to additionally reflect the complexity of the FJSSP benchmark instances as they quantify the degree of flexibility and the differences in processing times an instance is exposed to.
-[12]
+Moreover, the average number of operations per job $\nicefrac{{\texttt{N}}}{{n}}$, the machine flexibility $\beta$, and the duration variety $dv$ are provided.
+The latter two indicators are used to additionally reflect the complexity of the FJSSP benchmark instances as they quantify the degree of flexibility and the differences in processing times an instance is exposed to [32].
 The $\beta$-flexibility of a benchmark instance is determined as the ratio of the average amount of available machines per operation $m_{avg}$ and the total number of machines $m$ as
@@
 $10<\beta<$ 1, it is only partially-flexible and, consequently, harder to solve.
-The duration variety $dv$ is defined as the quotient of the number of unique values of the processing times $T_{ijk}$ within a FJSSP instance $d_{unique}$ and the total number of machine assignment options across all operations, i.e.,
-$\begin{split}dv=\frac{d_{unique}}{\sum_{i,j}\lvert\mathcal{M}_{i,j}\rvert}\end{split}$ (3)
+The duration variety $dv$ is defined as the quotient of the number of unique values of the processing times $T_{i,j,k}$ within a FJSSP instance $d_{unique}$ and the total number of machine assignment options across all operations, i.e.,
+$\begin{split}dv=\frac{d_{unique}}{\sum_{i,j}m_{i,j}}\end{split}$ (3)
 Low $dv$ values indicate either very similar processing times across all available machines or a large amount of assignment options, respectively.
 This corresponds to many distinct scheduling solutions that have the same level of fitness.
-Figure 3: Illustration of duration variety $dv$ and flexibility $\beta$ across all considered FJSSP instances.
+Figure 3: An overview over the benchmark instances on a subset of problem characteristics to show the distribution and similarities between instances.
+The instances are marked with respect to their original source.
 Figure 3 displays scatter plots of two characteristic features of all benchmark instances.
-This provides an first overview of the diversity of the individual instances and reveals where similarities coincide.
-Details on the individual characteristics of the considered benchmark sources can be found in the Appendix.
-This rough insight into the benchmark properties alone supports the demand for a targeted filtering of instance subgroups to get a better insight into the algorithmic working principles of developed solvers.
-2.3 Recent Work on Benchmarking Worker Flexibility in FJSSP
-In order to retrieve benchmark instances including worker flexibility, the instances are usually specifically constructed for the given use case.
-In most cases, the constructed instances are based on subsets of commonly used benchmark instances [14][15][16].
-In other cases, benchmark instances are randomly generated or constructed [16][17][18].
-While the construction process is sometimes explained ([19][14]), they are either adapted for a specific use case or only use a subset of the commonly used benchmark problems.
-The algorithm used to enable the conversion of all FJSSP benchmark instances was developed similarly to the algorithm used by [19].
-[18] tries to balance the makespan, maximum worker workload, and total workload of machines using an multi objective discrete teaching-learning based optimization approach.
+This provides a first overview of the diversity of the individual instances and reveals where similarities coincide.
+This illustration of benchmark properties supports the demand for targeted filtering of instance subgroups to get a better insight into the algorithmic working principles of developed solvers.
+Details on the particular characteristics of the considered benchmark sources can be found in the Github repository of the FJSSP(-W) benchmarking environment [41].
+The description of the respective environment is provided in Sec.
+4.
+3 Worker Flexibility
+Worker flexibility is a pivotal aspect of modern job scheduling, where the ability to adapt to varying worker skills and availability can significantly increase operational efficiency in production environments.
+This section addresses the Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W).
+It considers relevant benchmarking advances, provides a proper problem formulation, and discusses the design of the FJSSP-W instance set.
+3.1 Recent work on FJSSP-W benchmarking
+To obtain benchmark instances that take into account the flexibility of the workforce, the instances are usually designed specifically for the respective application.
+In most cases, the constructed instances are based on subsets of frequently used benchmark instances [43, 23, 24].
+Otherwise, benchmark instances are randomly generated or constructed [24, 44, 45].
+While the construction process is sometimes explained ([46, 43]), they are either adapted for a specific use case or only use a subset of the commonly used benchmark problems.
+The algorithm that enables the conversion of the FJSSP benchmark instances was developed similarly to the algorithm used by [47].
+However, to make the process easier to reproduce, the method to derive the durations based on the workers has been adapted to sample from a range around the original duration and default parameters for all other parameters (e.g.
+amount of workers) are given.
+In [45], the authors try to balance the makespan, maximum worker workload, and total workload of machines using a multi-objective discrete teaching-learning-based optimization approach.
 They use 30 newly constructed benchmark instances.
-[14] minimize makespan, labor cost and green production related factors using a non-linear integer programming model, as well as a non-dominated ensemble fitness ranking algorithm.
+[43] minimize the makespan, labor cost, and green production-related factors using a non-linear integer programming model, as well as a non-dominated ensemble fitness ranking algorithm.
 The 31 benchmark instances used, are constructed based on a subset of the commonly used FJSSP benchmark instances.
-[15] use a simulation based approach to optimize the FJSSP with worker flexibility and uncertain processing times with the minimum makespan and minimum standard deviation of makespan as objective.
-36 benchmark instances are used for evaluation, constructed based on a subset of the commonly used FJSSP instances.
-[16] compare a mixed-integer linear programming approach with NSGA-II in a distributed FJSSP with worker flexibility on a subset of the common benchmarks (50 instances) and randomly generated instances (8 instances).
-The objectives for their optimization are makespan minimization, balanced machine workload and workload of workers.
-[17] considers the JSSP with worker flexibility on newly generated benchmark instances, focusing on the maximum worker workload, minimizing the makespan and balancing the human energy expenditure.
-To evaluate their approach, 40 problem instances are generated.
-3 Representations of the Flexible Job Shop Scheduling Problem
-This section establishes the foundation for setting up the benchmarking environment by describing three different modeling approaches for the FJSSP.
-The first of these is a Mixed Integer Linear Programming (MILP) representation defined in Sec.
-3.1.
-It allows exact methods of mathematical programming to be applied.
-Furthermore, an implementation as Constraint Programming (CP) model is provided in Sec.
-3.2.
-Toolboxes for this type of problem representation are currently among the most sophisticated solving techniques for Production Scheduling tasks.
-Lastly, a convenient problem representation for Simulation-Based Optimization (SBO) algorithms, e.g.
-meta-heuristics, is presented in Sec.
-3.3.
+In [23], a simulation-based approach is used to optimize the FJSSP with worker flexibility and uncertain processing times.
+The objective considers minimizing both the makespan and the standard deviation of the makespan.
+There, 36 benchmark instances are used for evaluation, constructed based on a subset of the commonly used FJSSP instances.
+In [24], the authors compare a mixed-integer linear programming approach with NSGA-II in a distributed FJSSP with worker flexibility on a subset of the common benchmarks (50 instances) and randomly generated instances (8 instances).
+The objectives for their optimization are makespan minimization, balanced machine workload, and workload of workers.
+The paper [44] considers the JSSP with worker flexibility on newly generated benchmark instances, focusing on the maximum worker workload, minimizing the makespan, and balancing the human energy expenditure.
+For evaluation, 40 problem instances are generated.
+3.2 FJSSP-W problem formulation and instance characteristics
+The Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W) is an extension to the FJSSP which adds worker flexibility.
+This means that for every operation, an eligible worker also needs to be assigned to operate the machine to which the operation is assigned.
+The selected machine-worker combination further influences the processing time of the operation on the machine.
+Based on the notations [48], the FJSSP-W therefore introduces a set of $w$ workers $\mathcal{W}=\{W_{1},\dots,W_{w}\}$.
+Each operation $O_{i,j}$ needs to be assigned to a permissible machine $M_{k}\in\mathcal{M}_{i,j}\subset\mathcal{M}$ and a worker $W_{l}\in\mathcal{W}_{i,j,k}\subset\mathcal{W}$ the set of $w_{i,j,k}$ permissible workers to process the operation $O_{i,j}$ on the assigned machine $M_{k}$.
+The processing time of operation $O_{i,j}$ on machine $M_{k}\in M_{i,j}$ with worker $W_{l}\in W_{i,j,k}$ is given as $T_{i,j,k,l}$.
+Similar to the FJSSP, the FJSSP-W assumes that all operations, machines, and workers are operational and available at time 0.
+The objective for the FJSSP-W is to find those operation-machine-worker combination that minimize the makespan $C$.
+In contrast to the FJSSP case, the flexibility $\beta$ of an FJSSP-W benchmark instance needs to include more than only the operations and machines.
+All possible worker assignments for each operation-machine combination need to be taken into account.
+To this end, the flexibility as well as the duration variety provided for the FJSSP in Eqs.
+(2) and  (3) have to be adjusted to account for the additional worker requirements.
+Hence, for the FJSSP-W, the $\beta$-flexibility of a benchmark instance is determined as the ratio of the average amount of possible machine-worker combinations per operation $\omega_{avg}$ and the overall amount of unique machine-worker combinations available for the benchmark instance under consideration $\omega_{unique}$, i.e.,
+$\beta=\cfrac{\omega_{avg}}{\omega_{unique}}$ (4)
+To this end, the average amount of possible machine-worker combinations per operation $\omega_{avg}$ is calculated as
+$\omega_{avg}=\frac{1}{\texttt{N}}\sum_{i=1}^{n}\sum_{j=1}^{n_{i}}\sum_{k=1}^{m_{i,j}}w_{i,j,k}\quad$ (5)
+The flexibility value $\beta$ then indicates how many of the machines in the production environment can handle multiple operations [42].
+In the context of the FJSSP-W, it also indicates how many workers can handle several machines.
+If all existing assignment options are available for all operations, the flexibility gets $\beta=1$ and the FJSSP-W is called fully flexible.
+Otherwise, one obtains $10<\beta<$ 1.
+Such problems must be considered only partially-flexible, and consequently are harder to solve.
+The duration variety $dv$ is defined as the quotient of $d_{unique}$, the number of unique values of the processing times $T_{i,j,k,l}$ of the FJSSP-W instance, and the total number of operation-machine-worker combinations across all operations $\omega_{avg}\cdot\texttt{N}$, i.e.,
+$dv=\cfrac{d_{unique}}{\omega_{avg}\cdot\texttt{N}}$ (6)
+Low $dv$ values indicate very similar processing times across all possible combinations.
+This corresponds to many distinct candidate solutions that have the same level of fitness, leading to difficulties for optimization algorithms.
+Considering instance attributes, such as the total number of operations N, flexibility $\beta$, and duration variety $dv$, allows for a first intuition of the individual benchmark instance complexity.
+It also supports the performance analyses between different solvers, and enables a problem-dependent configuration of the solver-specific strategy parameters.
+3.3 Design of FJSSP-W instances
+The FJSSP-W considers that operations in the production process need workers to be present.
+Therefore, workers also need to be assigned to the different operations.
+In the FJSSP-W, it is considered that not all workers will work at the same speed, and not all workers are capable of performing every operation with ever (or any) of the available machines for the operation.
+To receive a set of benchmark problems to test and develop appropriate approaches, the given FJSSP benchmark instances are extended to contain the additional required information about the possible worker assignments.
+To that end, we provide an algorithm similar to [47] , which is capable of extending any FJSSP instance in the format of the gathered benchmark instances (Alg.
+1).
+The parameters used for the benchmark generation are the amount of workers (default: $1.5m$), the upper and lower bound range for the durations (default: $ub=1.1$, $lb=0.9$ of the original duration) to have a random fluctuation in the durations required by the workers of ±$\pm 10\%$, and optionally a random seed for reproducibility.
+Algorithm 1 Adding Worker Flexibility to common FJSSP Benchmark instances
+1:$w>0$, $lb\geq 0$, $ub\geq 0$, $path$
+2:$\mathbf{out}\leftarrow\{\}$
+3:$\mathbf{rows}\leftarrow ReadFile(path)$
+[…]
+5:for $i\leftarrow 1$ to $||\mathbf{rows}||$ do
+6:     $\mathbf{out}_{i}\leftarrow\{\}$, $idx\leftarrow 0$
+7:     $\mathbf{out}_{i}\cup\mathbf{rows}_{i,idx}$
+8:     for $j\leftarrow 1$ to $||\mathbf{rows}_{i}||$ do
+9:         $idx\leftarrow idx+1$
+10:         $m\leftarrow\mathbf{rows}_{i,idx}$
+11:         $\mathbf{out}_{i}\cup\{m\}$
+12:         for $k\leftarrow 1$ to $m$ do
+13:              $idx\leftarrow idx+1$
+14:              $\mathbf{out}_{i}\cup{\mathbf{rows}_{i,idx}}$
+15:              $workers\leftarrow Uniform(1,w)$
+16:              $\mathbf{out}_{i}\cup\{workers\}$
+17:              $\mathbf{selected}\leftarrow Sample(w,workers)$
+18:              $idx\leftarrow idx+1$
+19:              $d_{orig}\leftarrow\mathbf{rows}_{i,idx}$
+20:              for $l\leftarrow 1$ to $workers$ do
+21:                  $d_{work}\leftarrow Round(Uniform(d_{orig}*lb,d_{orig}*ub))$
+22:                  $\mathbf{out}_{i}\cup\{selected_{l},d_{work}\}$
+23:              end for
+24:         end for
+25:     end for
+26:end for
+After the initialization of the required parameters, in line 1 an empty list for saving the result benchmark is created.
+In line 2, the original benchmark file, located at $path$, is read and saved into a list, separated into the rows of the file.
+The base information about the benchmarks are added to the result, as a list representing a new row in the output, in line 3, keeping the same format as the original files, with the added information about the amount of workers available for the problem instance.
+Starting from line 4 and ranging to line 25, the rest of the rows of the original benchmark file are processed one by one.
+In line 5, a variable to keep track of the reading position in the current row ($idx$) is initialized, as well as an empty list, which is added to the output as a new row.
+The amount of operations needed for the job represented by the row is added to the new row in line 6.
+From line 7 to line 24, the current row from the original benchmark, according to the original format of the problem instances, is processed.
+To do so, $idx$ is raised by $1$ in row 8 to address the next operation that needs to be parsed.
+In line 9, the amount of machines available for the operation is read, and added to the output row in line 10.
+From line 11 to line 23, the different machine assignments and their respective durations for the current operation are parsed.
+To achieve that, line 12 raises $idx$ by $1$ again to advance the reading position of the input current row.
+Since $idx$ is now referencing to a machine ID in the original benchmark, this possible machine assignment is also added to the output row in line 13.
+In line 14, the amount of workers which are going to be able to perform the required operation on this specific machine is chosen uniformly from the interval [$1$, $w$].
+The amount of workers available for this machine assignment ($workers$) is added to the output row in line 15.
+In line 16, the available workers for the assignment are selected randomly, using a function named $Sample$, using the parameters $w$ and $workers$ to select the IDs of the workers in the interval [0, $w$).
+Note that generally the assigned IDs of machines and workers are converted to 0-based indexing for easier implementation as opposed to the original 1-based indexing.
+The subset of workers is stored in the list $\mathbf{selected}$.
+To advance the reading position again, $idx$ is raised by $1$ in line 17.
+$idx$ now points at the original duration of the machine assignment for the given operation, which is stored in $d_{orig}$ in line 18.
+Starting from line 19 and ending in line 22, the current operation-machine assignments are finalized by choosing the durations for the different worker assignments.
+To do this, the duration for each worker $d_{work}$ is uniformly chosen from the interval [$d_{orig}*lb$, $d_{orig}*ub$] and rounded to the closest integer in line 20.
+Line 21 adds the worker and the duration for the operation-machine-worker combination to the output row in $out$.
+After termination of this routine, the FJSSP-W benchmark is fully extended to include worker flexibility and it is then exported to a text file similar to the original FJSSP format described in Fig.
+2.
+Figure 4: Explanation of the FJSSP-W benchmark specifications based on an extended FJSSP test instance.
+As illustrated in Fig.
+4, the process in Alg.
+1 extends the first row of the original FJSSP instance description with a single number in third position.
+It adds one new value to specify the number of workers $n_{\omega}$ present in the production environment.
+Similar to the FJSSP format, the following rows represent the individual jobs $J_{i}$ $(i=1,\dots,n)$ to be completed by the production environment.
+For each of these jobs, the first value again gives the number of operations $n_{i}$ to be executed in job $J_{i}$.
+The second number is the number of machines $m_{i,1}$ available to process the first operation $O_{i,1}$ of job $J_{i}$.
+Now, in the FJSSP-W, each of these machines in $\mathcal{M}_{i,1}$ comes with a set of admissible workers $\mathcal{W}_{i,1,k}$.
+Consequently, the FJSSP format is modified in such a way that the machine-related processing times of the FJSSP instances are replaced by the admissible number of workers for the operation-machine combinations and corresponding processing times of those workers on that machine.
+In Fig.
+4, the modifications are highlighted by the black rectangles.
+The first value inside the rectangles indicates how many different workers can be assigned for the operation-machine combination.
+This leading number is followed by the corresponding pairs of worker IDs and associated processing times.
+Example given, the second job $J_{2}$ in the sample FJSSP-W instance shown in Fig.
+4 consists of $n_{2}=5$ operations.
+The first of these operations, i.e.
+$O_{2,1}$ can be executed on exactly one of the two available machines.
+It is machine $M_{1}$, which can be operated by exactly one of the three workers, namely the second worker $W_{2}$.
+Worker $W_{2}$ can carry out the operation $O_{2,1}$ on machine $M_{1}$ with processing time $T_{2,1,1,2}=58$.
+Analogously, the second operation $O_{2,2}$ of job $J_{2}$ can be performed on both machines ($M_{1}$ and $M_{2}$.
+On machine $M_{1}$ by one worker, namely worker $W_{3}$ with processing time $T_{2,2,1,3}=37$.
+In addition, the same operation $O_{2,2}$ can be performed on machine $M_{2}$ by two workers.
+Worker $W_{1}$ requires $T_{2,2,2,1}=30$ time units and worker $W_{3}$ has a processing time of $T_{2,2,2,3}=37$.
+4 Description of the Benchmarking Environment
+This section elaborates on the implementation details, usage options as well as performance assessment and visualization of the established benchmarking environment.
+4.1 Implementation details
+A Python implementation of the benchmarking environment is available on Github [41].
+Several APIs are offered by the environment, including the evaluation of FJSSP and FJSSP-W solutions and their comparison to the known results.
+The APIs include benchmark selection methods, evaluation methods as well as visualization methods for comparisons between solvers and (optionally) the provided best known results for the provided FJSSP and FJSSP-W instances.
+The repository is divided into a ”util” and ”instances” subsection.
+The ”util” section contains the provided APIs for comparison and evaluations, while the ”instances” subsection contains the problem instances, including the example FJSSP-W instances created by Alg.
+1, and information about the problem characteristics of the different problem instances.
+Furthermore examples for the usage of the APIs for the whole workflow are provided.
+For demonstration purposes the repository also contains a greedy search algorithm to showcase the results and how they can be compared to other solvers as well as the best known results.
+Additionally, the benchmarking environment contains the best known values for the FJSSP and FJSSP-W instances created with the same parameters than the instances used in this paper and provided by the benchmarking environment.
+The algorithm which turns FJSSP instances into FJSSP-W instances is also provided.
+4.2 Benchmark selection
+The benchmarking environment offers a filter option that allows benchmark instances to be selected according to the desired benchmark properties.
+Figure 5 shows the distribution of benchmark instances in relation to some of the filterable problem properties.
+There are several clusters of instances that are very similar in these characteristics.
+To efficiently test how well an algorithm applies to the more general problem, it may be desirable to first select only representatives of the clusters to save computation time and generate the most representative subset of instances possible to avoid potential biases in the results analysis.
+The filterable features include the number of operations N, the number of machines, the number of workers, the flexibility $\beta$, the duration variety $dv$, the number of orders $n$, and various other features (see [41]).
+Each characteristic can optionally be filtered with a predefined lower and upper limit for problem cases to be included in the experiments.
+This option is available for both the supplied FJSSP and FJSSP-W instances.
+4.3 Benchmark preparation
+Based on the provided filters (if any), the requested benchmarks can be loaded using built-in functions.
+The loading process returns all requested instances and provides several APIs to read their data.
+This includes the job sequence, which shows how many operations belong to each job, including their fixed position for the assignment vectors, the duration matrix $T$, the amount of operations N, machines $m$, worker $w$, and several other characteristics.
+In the case of the provided Greedy Solver, the duration matrix and the job sequence is the only required input to build its encoding and generate its solution.
+4.4 Benchmarking experiments
+The provided best results, both for the FJSSP and FJSSP-W, were retrieved from a combination of state-of-the-art solvers and a GA [32].
+The experiments to get the results were processed on the same test computer (Intel Core i7-6700 (3.40GHz) CPU with 16 GB RAM, Windows 10) with a time limit of 20 minutes.
+Note that for the FJSSP-W, only the best known results for the provided instances are given.
+If other FJSSP-W instances are generated, they are not comparable to the provided best known results.
+The visualization methods provided by the benchmarking environment can be used to visualize any metrics, though for comparison with the provided best known results the makespan is required as objective function.
+The best found lower bounds by the used solvers are also provided and can be compared.
+For the comparison of specific problem instances a progress plot can be created.
+This requires a record of the desired metrics over time in the form of a list.
+The list elements should contain tuples with the first entry being the timestamp of the record and the second being the recorded value.
+This data can be used by the visualizer to produce a comparative plot.
+Note that for solvers with a stochastic nature like metaheuristics, multiple experiments per problem instance should be performed to receive conclusive results.
+The minimum recommended amount of independent algorithm repetitions is 20.
+This represents a slight reduction from the usually recommended number of about 30 repetitions for statistically relevant results [49] and is due to the high computational effort for solver execution on the individual instances.
+To this end, it appears useful to speed up the duration of the experiments by carrying out the individual runs in parallel.
+However, the process of parallelization is left to the user of the benchmark environment and is not shown in the usage examples below, since parallelized approaches can not be tracked automatically.
+Yet, the best results and the average optimization results can be formatted into the required format and used with the provided visualization tools.
+(a)Refer to caption (b)Refer to caption
+Figure 5: Illustration of the distribution of benchmark instances with respect to their flexibility $\beta$, duration variety $dv$, and their total amount of operations N.
+Note that the calculations of $\beta$ and $dv$ differs between the FJSSP (c.f.
+Eqs.
+(2) and (3), and the FJSSP-W (c.f.
+Eqs.
+(4) and (6)), respectively.
+4.5 Performance evaluation
+The benchmark environment offers an evaluation function for the makespan to ensure the final solution is evaluated correctly.
+As input for the evaluation, the start times, the machine assignments and the worker assignments (in case of the FJSSP-W) are required in the fixed order provided by the benchmarking environment.
+A translation between the encoding used in [32] for both the FJSSP and the FJSSP-W is available.
+For comparisons to other solvers, the benchmark environment offers different visualizations of the results summarized over all used benchmark instances.
+The data of the evaluated solvers are included and best known results can be updated over time.
+Additionally, the benchmarking environment offers an API to calculate the MiniZinc score [50] for comparison.
+A demonstration can be found in the provided examples below.
+The example result plots are shown in Fig.
+6.
+For each compared scheduling algorithm, the illustration visualizes the proportion of problem instances that can be solved up to a given deviation from the best-known results on such instances.
+To this end, for each instance, the relative deviation, or gap, $\delta_{rel}$ of a solver’s final best solution $C_{fb}=C(\mathbf{y}_{fb})$ from the best-known result $C_{best}$ on a particular instance is calculated
+$\begin{split}\delta_{rel}=\frac{C_{fb}-C_{best}}{C_{best}}.\end{split}$ (7)
+All instances that could be solved within the deviation $\delta_{rel}$ are then counted.
+Finally, the ratio of instances solved within the gaps is then plotted against the respective $\delta_{rel}$ values.
+For multiple solvers, the algorithm performance is considered to be better the closer the curve runs along the vertical axis.
+In such cases, larger proportions of the instances are solved with comparatively smaller deviations from the best-known solution.
+While Fig.
+6(a) demonstrates a general overview of the performance differences among four exemplary solvers, Fig.
+6(b) zooms into the lower left part of Fig.
+6(a) to focus on more subtle performance differences between solvers.
+One observes that Solver A exhibits the best performance as it realizes the best-known solution on approximately $55\%$ of the instances.
+On the contrary, Solver C shows the poorest performance.
+Note that this representation is not suitable for drawing conclusions about the performance on individual instances but concentrates exclusively on the proportion of instances that could be solved to a given quality.
+(a)Refer to caption (b)Refer to caption
+Figure 6: An example plot using example data displaying the portion of instances that lie within a certain gap $\delta_{rel}$ from the best found solution.
+A gap of 1.0 corresponds to 100% distance (i.e.
+twice as high as the minimum) to the best solution.
+In (a), all of the data is displayed, while (b) shows an example with a more zoomed in perspective on the same data (limit on the x-axis as parameter to the visualizer code).
+As can be seen in Fig 6, it can happen that some performance lines do not reach $\delta_{r}el=n1.0$.
+This happens if not all solvers could find feasible solutions for all the included instances.
+The visualizer is implemented in a way that it can handle missing data for the case of Fig.
+6(a).
+MiniZinc Score and Statistical Significance Testing
+In addition to the performance plots in Fig.
+6, the benchmarking environment provides further performance indicators for comparing the tested solvers on all instances of both problem types (FJSSP or FJSSP-W).
+On the one hand, the suite offers the option of generating an overall score for the results of all benchmark instances, following the example of the annual MiniZinc Competition [51] of constraint programming solvers [52, 50].
+On the other hand, the rankings of the benchmarked solvers are computed and tested for statistically significant differences using a Friedman test as the omnibus test and, if applicable, the Nemenyi test for pairwise post-hoc investigations [53].
+Table 2: MiniZinc scores calculated for the example solvers. A higher score indicates superior performance.
+FJSSP-W solver Solver A Solver B Solver C Solver D
+MiniZinc score 866.0 590.0 460.0 496.0
+The so-called MiniZinc score is a summarized measure of the mutual performance of the compared solvers across all selected instances of a problem type.
+The score is calculated by comparing the results of each solver relative to the results of all other solvers on each instance.
+On every instance, a solver receives one point for every solver that achieves a worse solution quality or infeasible results.
+In case of an equal solution quality, the time required to achieve the result is also considered.
+In that case, the two compared solvers split the point weighted by their time to achieve the result.
+The maximum score of a solver is bound by the number of instances times the number of compared solvers minus 1, i.e.
+$MiniZinc\>score\>\leq\>\#instances\times\left(\#solvers-1\right).$ (8)
+For the example data used in the demonstration plots, the MiniZinc Score of the four solvers A to D is determined and displayed in Table 2.
+Considering the comparison of 4 solvers and the use of 402 instances the maximum score of a solver becomes 1206.
+For the purpose of a statistically sound statement about the significance of the performance differences, the compared solvers are ranked on the individual instances.
+The distributions of these ranks are then examined for significant differences using a Friedman test at the confidence level of $\alpha=5\%$.
+If the $p$-value returned by the Friedman test is below $\alpha$, there is an indication of significant differences.
+Consequently, a Nemenyi test is performed to make a pairwise comparison.
+The results of the statistical tests between all solvers are summarized in a Nemenyi diagram.
+Figure 7: A Nemenyi diagram based on the example data using 4 solvers for 402 problems.
+Figure 7 shows an example of a Nemenyi diagram for the four considered example solvers A to D together with the resulting $p$-value returned by the corresponding Friedman test.
+From left to right, the diagram orders the compared algorithms from inferior to superior ranks.
+It displays the average ranking of each solver as well as the critical distance $CD$ that needs to be exceeded to indicate a statistically significant performance difference.
+Below the range of admissible ranks, bold black lines connect the solvers that did not produce statistically significant differences.
+As can be seen in Fig.
+7, Solver A significantly outperforms the remaining three methods, while Solver C and Solver D produce statistically indistinguishable results.
+At the same time, Solver B and D do not show significant differences.
+Yet, Solver B exhibits a significantly better performance than Solver C.
+Single instance evaluation
+The plots shown in Fig.
+8 display the progress of a solver on a selected individual instance over time $t$ (measured in seconds).
+The example plots shows the relative deviation $\delta_{rel}$ (cf.
+Eq.
+(7) of the best solution found by each solver from the best solution observed by all considered solvers in Fig.
+8(a).
+Accordingly, the solver that first reaches $\delta_{rel}=0$ exhibits superior performance in terms of speed (i.e., Solver A in the provided example).
+Fig.
+8(b) displays the $\delta_{rel}$ dynamics towards a user-defined deviation (here 10%) from the best-known solution (here provided by Solver A).
+This target deviation is specified by the user and can be provided to the visualization tool as a parameter.
+Within the example data, all solvers reach the threshold distance to the best result in the given plots.
+However, if not all solvers reach the best result or get within the specified threshold (as used in 8(b)), not all lines might intersect the horizontal solid red line.
+This can also happen if a limit on the horizontal axis is imposed on the plot.
+Note that not all solvers start at $t=0$, since only feasible solutions are considered for the visualization of the dynamics.
+(a)Refer to caption (b)Refer to caption
+Figure 8: Illustration of the progress (7) of the example solvers on a specific problem.
+Part (a) displays the $\delta_{rel}$ dynamics relative to the best solution of all solvers over time $t$.
+Using a logarithmic scale on the y-axis, part (b) intends to show the $\delta_{rel}$ behavior until a user defined deviation (here 10%) from the best-known result (here Solver A) is reached.
+It indicates the time of inferior solvers to approach a comparable solution quality.
+The user defined target deviation is represented by the horizontal solid red line.
+5 Example Experiments
+The benchmarking environment is demonstrated by comparing different solvers on the FJSSP-W benchmarks.A complete example (from filtering and loading the data to visualizing the result) comparing the provided Greedy solver with the best known results is provided with the environment.
+The implementation of the Greedy solver is intended to facilitate the application of the benchmark environments for users.
+Its results can be regarded as absolute baseline results, which should in any case be surpassed on the individual instances.
+5.1 Experimental Settings
+For the FJSSP-W, the solvers were tasked with solving all 402 benchmarks.
+The hardware used for the experiments is a Intel Core i7-6700 (Quad-Core, 3.40 GHz) CPU with 16 GB RAM.
+All solvers were given a maximum of 1200 seconds (20 minutes) for each benchmark instance.
+The respective benchmark instances, including all final results, are available online at [54].
+5.2 Applied Solvers
+In order to test the benchmarking environment to compare different solvers for the FJSSP-W, a Genetic Algorithm (GA) [32], a Mixed-Integer Linear Programming solver (MILP), as well as a Constraint Programming solver (CP) are compared to a simple Greedy algorithm as baseline.
+The MILP and CP solver used for the comparison are both provided by the CPLEX library developed by IBM [29].The Greedy algorithm considers the next operation of the sequence of every job as next operation to insert into the schedule.
+The operation with the smallest machine-worker combination is chosen.
+If multiple operations offer the same fastest time, the operation is chosen randomly between the best options.
+5.3 Results
+(a)Refer to caption
+(b)Refer to caption
+(c)  Solver CP MILP GA Greedy Score 1160.5 273.0 778.5 200.0
+Figure 9: The $\delta_{rel}$ for the FJSSP-W on benchmark instances for the solvers used for the comparison of the approaches is shown in (a).
+In (b), the Nemenyi diagram for the used solvers is depicted showing the average rankings of each solver.
+The table in (c) shows the MiniZinc score achieved by the compared solvers.
+Figure 9(a) shows that the CP as well as the GA approach achieve values of $\delta_{rel}<0.25$ on more than $80\%$ of the 402 benchmark instances.
+Yet, neither the MILP solver nor the Greedy approach get anywhere near a similar performance.
+While the greedy approach would reach $100\%$ eventually, the MILP approach could not solve all problem instances, due to the large amount of variables required and the imposed time limit.
+Furthermore, hardware limits for large problem instances cause the MILP approach to run out of memory due to the high number of created constraints.
+Figure 9(b) shows that the CP solver significantly outperforms the other approaches.
+The second placed GA approach also significantly outperforms the MILP and Greedy solvers that do not show significantly different performances over all instances.
+The MiniZinc score shown in Fig.
+9(c) also support the results shown in Fig.
+9(b), with the CPs score being considerably higher than the other three solvers, while the GAs score is still significantly higher than the MILP and Greedy score.
+The MILP and Greedy approach achieve a similar score, with the MILP approach ultimately scoring higher than the Greedy algorithm.
+6 Conclusion
+This paper introduced a comprehensive benchmarking environment for the Flexible Job Shop Scheduling Problem (FJSSP) and its extension, the FJSSP with Worker Flexibility (FJSSP-W).
+It also offers a method to convert FJSSP instances into FJSSP-W instances.
+The benchmarking environment enables the comparison of solver results against the best-known solutions.
+Additionally, it provides tools to filter problem instances based on specific characteristics and to visualize the (relative) performance of various solvers w.r.t.
+the best-known results of state-of-the-art solvers.
+Benchmarking optimization solvers is a critical process that requires a consistent and systematic approach to ensure reliable and insightful algorithm quality assessments.
+To achieve this, several recommendations [5, 9, 10] need to be followed and adopted for use of the introduced FJSSP(-W) benchmark environment.
+For instance, it is essential to select a most representative set of problem instances that reflect the diversity tied to the benchmarking objectives.
+To this end, the proposed FJSSP(-W) environment is based on a set of 402 commonly accepted and diverse FJSSP instances.
+These are systematically extended to account for worker flexibility (FJSSP-W).
+Although the benchmarking environment includes the code for the independent creation of corresponding FJSSP-W instances, it is strongly recommended to use the fixed and provided 402 FJSSP-W instances of the repository in order to ensure the comparability and replicability of a benchmark suite.
+This is the only way to generate a comprehensive picture of the performance of a solver in relation to the best known results provided and to future developments.
+Further, the creation of additional new instances can be useful in order to investigate certain characteristics in more detail and to expand the space of FJSSP-W test instances in a meaningful way.
+However, such adaptations should be clearly indicated in the interests of honest scientific work.
+The use of the filter to select subsets of the 402 instances is intended in particular for the targeted algorithmic development of operators that focus on specific aspects of the problem.
+By avoiding the need to evaluate all 402 instances at all times, valuable development time can potentially be saved.
+Moreover, choosing appropriate performance metrics is important.
+To start with, this version of the environment is only designed to measure the makespan.
+This is a common and useful performance measure for both the FJSSP and the FJSSP-W.
+In future updates, the environment will be expanded to include further performance measures.
+As a computing time limit, we suggest a duration of 20 minutes per algorithm run.
+This time limit was also used for the present reference results.
+To grade the computing environment and understand the solver’s performance in practical terms, users are expected to report their hardware specifications and the computational resources used.
+To ensure that the experiments can be replicated by others, it is necessary to provide detailed descriptions of the setup and parameters.
+This applies for the parameters of the benchmarking environment (it is recommended to keep the default parameters in Sec.
+4) as well as for the strategy parameters of the solvers.
+Whether using default settings or custom configurations, transparency in describing the algorithms and experimental conditions is key.
+This allows other researchers to reproduce the experiments and verify the results.
+Since the benchmark environment is clearly aimed at heuristic scheduling procedures and evolutionary algorithms in particular, it is necessary to take the stochasticity of these algorithms into account.
+This is another important aspect to evaluate the stability of the solver and to understand the reliability and consistency of the results.
+To this end, developers of such solvers are demanded to conduct repeated experiments and to analyze the fluctuations in the results.
+The present paper recommends a minimum amount of $20$ repetitions on each instance to appropriately take into account random effects.
+The systematic analysis of the results is essential to gain meaningful insights.
+In order to interpret the results and determine the significance of the differences between the solvers, a statistical analysis should be carried out.
+In this way, it is possible to determine which solvers perform better under certain conditions and a solid basis for comparison is created.
+To ensure this, the present benchmarking environment offers a ranking mechanism that statistically evaluates the performance of several solvers based on the Friedman and Nemenyi test and summarizes it in a corresponding diagram.
+The solver results need to be clearly and intuitively presented to visually support the understanding of the performance differences and to communicate experimental findings effectively.
+With the representation of the overall performance in Fig.
+6 and the instance-related makespan dynamics in Fig.
+8, the environment simplifies the visual judgment of the performance results on this broad set of benchmark instances.
+Transparency in reporting benchmarking findings is essential.
+All relevant code and benchmark data need to be made accessible to the research community.
+This facilitates subsequent studies and comparisons, contributing to the collective knowledge in the field.
+Further, the related Github repository ensures continuing support to regularly update the benchmarking processes in consultation with the research community.
+This helps in refining benchmarking practices and addressing emerging challenges, leading to more accurate and reliable comparisons over time.
+Following such guidelines establishes a consistent and systematic approach to benchmarking FJSSP(-W) optimization solvers which will drive progress in this field of research.
+7 Outlook
+In future work for the FJSSP and FJSSP-W, additional evaluation metrics (e.g., schedule tardiness or robustness) can be added to the benchmarking environment.
+In addition, models and simulations for added uncertainty to the provided problem instances can also be included.
+Uncertainties for the FJSSP include uncertain processing times, machine breakdowns, dynamic job arrivals.
+Alongside with models and simulations for these types of uncertainties, performance metrics to evaluate the robustness of the derived schedules can also be added to further support the development solver for closer to real-world use cases.
+For easier analysis of the results, additional visualizations might also be added.
+Considering the different characteristics of the problem instances, an analysis tool indicating which solver performed best on which benchmark characteristics in direct comparison might also be a valuable extension.
+Acknowledgements
+The financial support by the Austrian Federal Ministry of Labour and Economy, the National Foundation for Research, Technology and Development and the Christian Doppler Research Association is gratefully acknowledged.
+Supplementary Material
+By describing three different problem representations for the FJSSP(-W) instances, this supplementary section establishes the foundation to utilize the benchmarking environment for different types of solvers.
+This is intended to facilitate the comparison of algorithmic results with solvers from other optimization areas.
+In so doing, we do not claim to be universally valid and complete, but merely provide suggestions.
+The use of own and possibly more efficient problem formulations is not excluded, taking into account the benchmarking principles of reproducibility and comparability.
+The following sections describe three types of problem formulations.
+The formulations are provided for the FJSSP instances, and for the FJSSP-W instances, respectively.
+In Sec.
+A, a Mixed Integer Linear Programming (MILP) representation is presented that allows for the application of exact Mathematical Programming solvers like Gurobi.
+Further, an implementation as Constraint Programming (CP) model is provided in Sec.
+B.
+Toolboxes (e.g.
+CPLEX) for this type of problem representation are currently among the most sophisticated solving techniques for Production Scheduling tasks.
+Finally, a convenient problem representation for Simulation-Based Optimization (SBO) algorithms (e.g.
+meta-heuristics, like GA) is presented in Sec C.
 It includes the definition of a precise schedule encoding and the design of the production system as an functional simulation environment.
-3.1 Mixed-Integer Linear Programming Formulation
-The Mixed Integer Linear Program (MILP) formulation is a proven model for the FJSSP to address these problem class using mathematical optimization methods, i.e.
+Appendix A Mixed-Integer Linear Programming Formulation
+The Mixed Integer Linear Program (MILP) formulation is a common model to approach the FJSSP(-W) problem class using mathematical optimization methods, i.e.
 branch-and-bound or cutting-plane methods.
-In this section, the MILP is systematically formulated describing the decision variables, constraints and objective functions in detail.
-Accordingly, the decision variables are defined in the following way:
+In this section, the MILP representation is systematically formulated describing the decision variables, constraints and objective functions in detail.
+All necessary parameters are first specified for the FJSSP before an extension for the FJSSP-W is provided.
+A.1 Decision Variables of the FJSSP
+In a first step, the decision variables can be defined in the following way:
 Notation Description of decision variables Domain
@@
 Afterwards, address a typical optimization objective associated constraint formulations.
-3.1.1 General restrictions in FJSSP
+A.2 General Restrictions in FJSSP
 Assignment constraints
@@
 (10), (11), (12), (14), (15), and (16).
-3.1.2 Objective function and associated constraints
+A.3 Objective function and associated constraints for the FJSSP
 Makespan minimization
@@
 $C_{max}\geq C_{i,n_{i}}\qquad\text{ for }1\leq i\leq n.$ (18)
-3.2 Constraint Programming Formulation
-The FJSSP can be easily formulated as a discrete finite constraint optimization problem(COP).
-The two COP solvers that we take into account in this comparison implement variable types, like interval variables and constraints like no-overlap or end-before-start, that are tailored to deal with scheduling problems.
-In this comparison OR-Tools (ORT) an open-source solver developed by Google and CP-Optimizer (CPO) a constraint optimization solver provided as a part of the IBM ILOG CPLEX Optimization Studio are used.
-The problem formulation for both solvers uses the same subscripts, variables, and index sets as the MILP formulation (see 3.1).
-3.2.1 Variables
+A.4 Extended MILP Formulation for the FJSSP-W
+An obvious generalization of the MILP formulation for the FJSSP (LABEL:sec:fjssp-milp) is used.
+The worker flexibility is modeled by an additional set of constraints that follows the same ideas as the sequencing constraints on machine level in LABEL:sec:fjssp-milp.
+The following description explains in detail only the additional notations and constraints of the MILP formulation for the FJSSP-W.
+A.4.1 Indices and Variables
+To denote the workers the subscripts $s,s^{\prime}$ with $1\leq s,s^{\prime}\leq w$ and $w$ the number of workers are used.
+The processing time of operation $O_{i,j}$ on machine $M_{k}$ also depends on worker $W_{s}$ and is denoted by $T_{i,j,k,s}$
+The decision variables $Y_{i,j,k}$ in LABEL:sec:fjssp-milp have to be changed to $Y_{i,j,k,s}$ which are 1, if operation $O_{i,j}$ is processed on machine $M_{k}$ by worker $W_{s}$ and 0 else.
+Furthermore an additional type of decision variables is necessary: $U_{i,j,i^{\prime},j^{\prime}}$, which is set to 1 if Operation $O_{i,j}$ is processed after operation $O_{i^{\prime},j^{\prime}}$ by the same worker.
+To describe the admissible workers for processing operation $O_{i,j}$ on machine $M_{k}$ the index set
+$W_{i,j,k}=\{s:\text{ such that $O_{i,j}$ can be processed on $M_{k}$ by worker$W_{s}$}\}$
+is used.
+A.4.2 Constraints
+Assignment constraints
+These constraints guarantee that all operations $O_{i,j}$ are assigned to exactly one machine and one worker.
+$\sum\nolimits_{k\in M_{i,j}}\sum\nolimits_{s\in W_{i,j,k}}Y_{i,j,k,s}=1\text{ for }1\leq i\leq n,1\leq j\leq n_{i}$
+Sequencing constraints on job level
+$C_{i,1}\geq\sum\nolimits_{k\in M_{i,1}}\sum\nolimits_{s\in W_{i,1,k}}Y_{i,1,k,s}T_{i,1,k,s}\text{ for }1\leq i\leq n$
+$C_{i,j}\geq C_{i,j-1}+\sum\nolimits_{k\in M_{i,j}}\sum\nolimits_{s\in W_{i,j,k}}Y_{i,j,k,s}T_{i,j,k,s}\text{ for }1\leq i\leq n,1<j\leq n_{i}$
+Sequencing constraints on machine level
+$C_{i,j}\geq C_{i^{\prime},j^{\prime}}+T_{i,j,k,s}-L(3-X_{i,j,i^{\prime},j^{\prime}}-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k,s^{\prime}})$
+$C_{i^{\prime},j^{\prime}}\geq C_{i,j}+T_{i^{\prime},j^{\prime},k,s^{\prime}}-L(X_{i,j,i^{\prime},j^{\prime}}+2-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k,s^{\prime}})$
+Both constraints have to hold for operations $O_{i,j}$ with $1\leq i<n$ and $1\leq j\leq n_{i}$ and operation $O_{i^{\prime},j^{\prime}}$ with $i^{\prime}>i$ and $1\leq j^{\prime}\leq n_{i^{\prime}}$, for machines $M_{k}$ with $k\in M_{i,j}\cap M_{i^{\prime},j^{\prime}}$ and for workers $W_{s}$ ($s\in W_{i,j,k}$) and $W_{s^{\prime}}$ ($s^{\prime}\in W_{i^{\prime},j^{\prime},k}$).
+Sequencing constraints on worker level
+$C_{i,j}\geq C_{i^{\prime},j^{\prime}}+T_{i,j,k,s}-L(3-U_{i,j,i^{\prime},j^{\prime}}-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k^{\prime},s})$
+$C_{i^{\prime},j^{\prime}}\geq C_{i,j}+T_{i^{\prime},j^{\prime},k^{\prime},s}-L(U_{i,j,i^{\prime},j^{\prime}}+2-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k^{\prime},s})$
+As in A.4.2 these two sets of constraints have to hold for operations $O_{i,j}$ with $1\leq i<n$ and $1\leq j\leq n_{i}$ and operation $O_{i^{\prime},j^{\prime}}$ with $i^{\prime}>i$ and $1\leq j^{\prime}\leq n_{i^{\prime}}$.
+But contrary to A.4.2 they have to hold for machines $M_{k}$ ($k\in M_{i,j}$) and $M_{k^{\prime}}$ ($k^{\prime}\in M_{i^{\prime},j^{\prime}}$) and for workers $W_{s}$ with $s\in W_{i,j,k}\cap W_{i^{\prime},j^{\prime},k^{\prime}}$.
+For the large number $L$, one can use
+$L=\sum\nolimits_{i=1}^{n}\sum\nolimits_{j=1}^{n_{i}}\max\nolimits_{1\leq k\leq m;1\leq s\leq w}(T_{i,j,k,s})$
+The idea of these constraints is the same as in Eq. (14) and (15) explained there in detail. Finally makespan and variable type constraint are the same as in A.3.
+Appendix B Constraint Programming Formulation
+The FJSSP can be easily formulated as a discrete finite single-objective Constraint Optimization Problem (COP).
+The two Constraint Programming (CP) solvers that we take into account in this comparison implement variable types, like interval variables and constraints like no-overlap or end-before-start, that are tailored to deal with scheduling problems of general form.
+The investigations included the solvers OR-Tools (ORT), i.e.
+an open-source solver developed by Google, and CP-Optimizer (CPO), a constraint optimization solver provided as a part of the IBM ILOG CPLEX Optimization Studio.
+The problem formulation for both solvers uses the same subscripts, variables, and index sets as the MILP formulation (see A).
+B.1 Variables
 The parameters representing the data are $T_{i,j,k}$ the processing time of operation $O_{i,j}$ on machine $k$.
@@
 The ORT solver additionally uses a boolean variable $a\_presence(i,j,k)$ which indicates whether the alternative interval $a\_I(i,j,k)$ is chosen.
-3.2.2 Objective
+B.2 Objective
 The objektive, i.e. minimizing the makespan, is formulated as
 $\max_{1\leq i\leq n}(end\_of(I_{i,n_{i}}))\to\min$
-3.2.3 Constraints
+B.3 Constraints
 Assignment constraints
@@
 $no\_overlap\left(a\_I(i,j,k):\mbox{ for }1\leq i\leq n,1\leq j\leq n_{i}\mbox{if }k\in M_{i,j}\right)$
-3.3 Simulation-Based Optimization Attempt
+B.4 Extending the CP Formulation for the FJSSP-W
+The same as for the MILP formulation holds also for the CP formulation: The FJSSP-W is an obvious generalization of the FJSSP.
+In the following the formulation for CPO solver is presented.
+The necessary changes for using the ORT solver are obvious when the two different formulations for the FJSSP are kept in mind.
+Again, each operation is represented by an interval variable $I(i,j)$ with starting time $start(i,j)$, ending time $end(i,j)$ and duration $duration(i,j)$.
+For the FJSSP-W an alternative interval variable $a\_I(i,j,k,s)$ has to be defined for each admissible machine-worker combination.
+The assignment of an operation $O_{j,k}$ to exactly one machine-worker combination is done by linking the interval variable $I(j,k)$ to an alternative interval variable $a\_I(i,j,k,s)$.
+This can be done by the $alternative$ constraints
+$alternative\left(I(i,j),[a\_I(i,j,k,s):\mbox{ for }k\in M_{i,j},s\in W_{i,j,k}]\right)$
+for $1\leq i\leq n$ and $1\leq j\leq n_{i}$.
+The sequencing constraints on job level are the same as in LABEL:sec:fjssp-cp:
+$end\_before\_start(I(i,j-1),I(i,j))\mbox{ for }1\leq i\leq n\mbox{ and }2\leq j\leq n_{i}$
+and the sequencing constraints on machine level are very similar to LABEL:sec:fjssp-cp. For $1\leq k\leq m$ the following constraints are added:
+$no\_overlap\left(a\_I(i,j,k,s):\mbox{ for }1\leq i\leq n,1\leq j\leq n_{i}\mbox{ if }k\in M_{i,j}\mbox{ and }s\in W_{i,j,k}\right)$
+Additionally the sequencing on worker level has to be guaranteed by using another set of $no\_overlap$ constraints. For $1\leq s\leq w$:
+$no\_overlap\left(a\_I(i,j,k,s):\mbox{ for }1\leq i\leq n,1\leq j\leq n_{i},k\in M_{i,j}\mbox{ if }s\in W_{i,j,k}\right)$
+Appendix C Simulation-Based Optimization Attempt
 Multiple problem encodings can be found in the literature, the most common one being [46, 32].
@@
 $C_{\max}=\max_{\tiny\begin{array}[]{c}1\leq i\leq n\\ 1\leq j\leq n_{i}\end{array}}(t_{i,j}+d_{i,j})$ (23)
-4 Expansion of the FJSSP for Worker Flexibility
-The Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W) is an extension to the FJSSP which adds worker flexibility.
-This means that for every operation, an eligible worker also needs to be assigned to operate the machine to which the operation is assigned.
-The selected machine-worker combination further influences the processing time of the operation on the machine.
-In addition to the FJSSP, the FJSSP-W therefore introduces a set of workers $\mathcal{W}=\{W_{1},\dots,W_{o}\}$.
-Each operation $O_{i,j}$ needs to be assigned to a permissible machine $M_{k}\in\mathcal{M}_{i,j}\subset\mathcal{M}$ and a worker $W_{s}\in\mathcal{W}_{i,j,k}\subset\mathcal{W}$ the set of permissible workers to process the operation $O_{i,j}$ on the assigned machine $M_{k}$.
-Similar to the FJSSP, the FJSSP-W assumes that all operations, machines, and workers are operational and available at time 0.
-The objective for the FJSSP-W is to find those operation-machine-worker combination that minimize the makespan $C$.
-As opposed to the FJSSP, the flexibility $\beta$ of a benchmark instance needs to take more into account than only the machines.
-All possible assignment combination for every operation need to be taken into account.
-This is the same as for the FJSSP, but includes the possible worker assignments for each operation-machine combination as well.
-The average amount of assignment options $\omega_{avg}$ is determined by Eq.
-(20).
-$\omega_{all}=\sum_{i=1}^{\texttt{N}}(|\mathcal{M}_{i}|+\sum_{j=1}^{|\mathcal{M}_{i}|}|\mathcal{W}_{\mathcal{M}_{i},j}|)$ (19)
-$\omega_{avg}=\frac{1}{\texttt{N}}\sum_{i=1}^{n}\sum_{j=1}^{n_{i}}\sum_{k=1}^{m}|\mathcal{W}_{ijk}|\quad$ (20)
-Which leads to $\beta$ defined as Eq. (21), with $\omega_{uni}$ being the overall amount of unique machine-worker combinations available for the benchmark instance.
-$\beta=\cfrac{\omega_{avg}}{\omega_{uni}}$ (21)
-In the same vein, the definition for $dv$ for FJSSP-W needs to be rewritten as Eq. (22)
-$\begin{split}dv=\cfrac{d_{unique}}{\sum_{i=1}^{\texttt{N}}(|\mathcal{M}_{i}|+\sum_{j=1}^{|\mathcal{M}_{i}|}|\mathcal{W}_{\mathcal{M}_{i},j}|)}\end{split}$ (22)
-4.1 Design Proposal of Worker Flexibility Integration
-The FJSSP-W considers that operations in the production process need workers to be present.
-Therefore, workers also need to be assigned to the different operations.
-In the FJSSP-W, it is considered that not all workers will work at the same speed, and not all workers are capable of performing every operation with ever (or any) of the available machines for the operation.
-To receive a set of benchmark problems to test and develop appropriate approaches, the given FJSSP benchmark instances are extended to contain the additional required information about the possible worker assignments.
-To that end, we provide an algorithm similar to [20] , which is capable of extending any FJSSP instance in the format of the gathered benchmark instances (Alg.
-1).
-The parameters used for the benchmark generation are the amount of workers (default: $1.5*m$), the upper and lower bound range for the durations (default: $ub=1.1$, $lb=0.9$ of the original duration) to have a random fluctuation in the durations required by the workers of ±$\pm 10\%$, and optionally a random seed for reproducibility.
-After the initialization of the required parameters, in line 1 an empty list for saving the result benchmark is created.
-In line 2, the original benchmark file, located at $path$, is read and saved into a list, separated into the rows of the file.
-The base information about the benchmarks are added to the result, as a list representing a new row in the output, in line 3, keeping the same format as the original files, with the added information about the amount of workers available for the problem instance.
-Starting from line 4 and ranging to line 25, the rest of the rows of the original benchmark file are processed one by one.
-In line 5, a variable to keep track of the reading position in the current row ($idx$) is initialized, as well as an empty list, which is added to the output as a new row.
-The amount of operations needed for the job represented by the row is added to the new row in line 6.
-From line 7 to line 24, the current row from the original benchmark, according to the original format of the problem instances, is processed.
-To do so, $idx$ is raised by $1$ in row 8 to address the next operation that needs to be parsed.
-In line 9, the amount of machines available for the operation is read, and added to the output row in line 10.
-From line 11 to line 23, the different machine assignments and their respective durations for the current operation are parsed.
-To achieve that, line 12 raises $idx$ by $1$ again to advance the reading position of the input current row.
-Since $idx$ is now referencing to a machine ID in the original benchmark, this possible machine assignment is also added to the output row in line 13.
-In line 14, the amount of workers which are going to be able to perform the required operation on this specific machine is chosen uniformly from the interval [$1$, $n_{workers}$].
-The amount of workers available for this machine assignment ($workers$) is added to the output row in line 15.
-In line 16, the available workers for the assignment are selected randomly, using a function named $Sample$, using the parameters $n_{workers}$ and $workers$ to select the IDs of the workers in the interval [0, $n_{workers}$).
-Note that generally the assigned IDs of machines and workers are converted to 0-based indexing for easier implementation as opposed to the original 1-based indexing.
-The subset of workers is stored in the list $\mathbf{selected}$.
-To advance the reading position again, $idx$ is raised by $1$ in line 17.
-$idx$ now points at the original duration of the machine assignment for the given operation, which is stored in $d_{orig}$ in line 18.
-Starting from line 19 and ending in line 22, the current operation-machine assignments are finalized by choosing the durations for the different worker assignments.
-To do this, the duration for each worker $d_{work}$ is uniformly chosen from the interval [$d_{orig}*lb$, $d_{orig}*ub$] and rounded to the closest integer in line 20.
-Line 21 adds the worker and the duration for the operation-machine-worker combination to the output row in $out$.
-Algorithm 1 Adding Worker Flexibility to common FJSSP Benchmark instances
-1:$n_{workers}>0$, $lb\geq 0$, $ub\geq 0$, $path$
-2:$\mathbf{out}\leftarrow\{\}$
-3:$\mathbf{rows}\leftarrow ReadFile(path)$
-[…]
-5:for $i\leftarrow 1$ to $||\mathbf{rows}||$ do
-6:     $\mathbf{out}_{i}\leftarrow\{\}$, $idx\leftarrow 0$
-7:     $\mathbf{out}_{i}\cup\mathbf{rows}_{i,idx}$
-8:     for $j\leftarrow 1$ to $||\mathbf{rows}_{i}||$ do
-9:         $idx\leftarrow idx+1$
-10:         $m\leftarrow\mathbf{rows}_{i,idx}$
-11:         $\mathbf{out}_{i}\cup\{m\}$
-12:         for $k\leftarrow 1$ to $m$ do
-13:              $idx\leftarrow idx+1$
-14:              $\mathbf{out}_{i}\cup{\mathbf{rows}_{i,idx}}$
-15:              $workers\leftarrow Uniform(1,n_{workers})$
-16:              $\mathbf{out}_{i}\cup\{workers\}$
-17:              $\mathbf{selected}\leftarrow Sample(n_{workers},workers$)
-18:              $idx\leftarrow idx+1$
-19:              $d_{orig}\leftarrow\mathbf{rows}_{i,idx}$
-20:              for $l\leftarrow 1$ to $workers$ do
-21:                  $d_{work}\leftarrow Round(Uniform(d_{orig}*lb,d_{orig}*ub))$
-22:                  $\mathbf{out}_{i}\cup\{selected_{l},d_{work}\}$
-23:              end for
-24:         end for
-25:     end for
-26:end for
-After this routine is finished, the benchmark is extended with worker flexibility and kept similar to the original format.
-The original format can be seen in an arbitrarily constructed example in Listing 2.
-12 3
-23 2 0 10 2 15 3 0 10 1 5 2 10 2 0 10 2 15
-32 1 1 15 2 1 10 2 15
-Listing 2: Original Benchmark example
-Row 1 of Listing 2 contains the information that the problem instance consists of 2 jobs and 3 available machines.
-In row 2 it is specified that Job 1 consists of 3 operations.
-Operation 1 of Job 1 can be processed on 2 machines (Machine 0 with processing time 10 and Machine 1 with processing time 15).
-Operation 2 of Job 1 can be processed on 3 machines.
-Either Machine 0 with processing time 10, Machine 1 with processing time 5 or Machine 2 with processing time 10.
-Operation 3 of Job 1 has 2 machines available (Machine 0 with processing time 10 and Machine 2 with processing time 15).
-Job 2 requires 2 operations.
-Operation 1 of Job 2 can be processed on only 1 machine (Machine 1 with processing time 15), while Operation 2 of Job 2 can be processed on 2 machines (Machine 1 with processing time 10 and Machine 2 with processing time 15).
-In the new format of the benchmarks, line 1 contains the general information about the instance (amount of jobs, amount of machines, amount of workers), followed by a row by row description of every job that needs to be scheduled.
-Every row starts with the number of operations required for the job to be finished.
-Following this first row entry, every operation starts with the amount of machines required, followed by the entries for every machine.
-Every entry for the machine starts with the machine id, followed by the amount of workers available to process the operation on this specific machine.
-After that, for every worker available on the machine, an entry consisting of the worker id, followed by the duration of the operation-machine-worker combination is listed in the row.
-An arbitrarily constructed example for this format can be seen in Listing 3. REPLACE:
-1 2 3 4
-2 3 2 0 2 1 10 3 5 2 1 0 15 3 0 1 2 10 1 2 0 10 3 5 2 2 2 5 3 15 2 0 2 2 15 3 10 2 2 0 15 1 5
-3 2 1 1 3 1 10 2 5 3 10 2 1 1 2 10 2 2 0 10 3 15
-Listing 3: Benchmark with worker flexibility example
-This example means that, according to row 1, there are 2 jobs, 3 machines and 4 workers in this production environment.
-Job 1 (row 2) consists of 3 operations.
-Operation 1 can be processed on 2 machines.
-The first of these machines is Machine 0, which has 2 workers available to perform this operation.
-Worker 1 with a processing time 10 and Worker 3 with a processing time of 5.
-The second machine to process Operation 1 of Job 1 is Machine 2.
-Machine 2 has 1 worker available (Worker 0) with a processing time 15.
-Operation 2 of Job 1 has 3 machines available.
-Machine 0, with 1 worker (Worker 2) with processing time 10, Machine 1 with 2 workers (Worker 0 with processing time 10 and Worker 3 with processing time 5), and Machine 2 with 2 workers (Worker 2 with processing time 5 and Worker 3 with processing time 15).
-Operation 3 of Job 1 can be processed on 2 machines.
-Machine 0, which has 2 possible worker assignments (Worker 2 with processing time 15 and Worker 3 with processing time 10), and Machine 2.
-Machine 2 also has 2 possible workers (Worker 0 with processing time 15 and Worker 1 with processing time 5).
-Job 2 (row 3), requires 2 operations to be processed.
-Operation 1 of Job 2 has 1 possible machine (Machine 1) with 3 possible workers.
-Worker 1 with processing time 10, Worker 2 with processing time 5 and Worker 3 with processing time 10.
-Operation 2 of Job 2 can be assigned to 2 machines.
-Either Machine 1, which has 1 possible worker (Worker 2 with processing time 10), or Machine 2, with the possibility to assign the operation to 2 different workers (Worker 0 with processing time 10 or Worker 3 with processing time 15).
-4.2 Mixed-Integer Linear Programming Formulation
-An obvious generalization of the MILP formulation for the FJSSP (3.1) is used.
-The worker flexibility is modeled by an additional set of constraints that follows the same ideas as the sequencing constraints on machine level in 3.1.
-The following description explains in detail only the additional notations and constraints of the MILP formulation for the FJSSP-W.
-4.2.1 Indices and Variables
-To denote the workers the subscripts $s,s^{\prime}$ with $1\leq s,s^{\prime}\leq w$ and $w$ the number of workers are used.
-The processing time of operation $O_{i,j}$ on machine $M_{k}$ also depends on worker $W_{s}$ and is denoted by $T_{i,j,k,s}$
-The decision variables $Y_{i,j,k}$ in 3.1 have to be changed to $Y_{i,j,k,s}$ which are 1, if operation $O_{i,j}$ is processed on machine $M_{k}$ by worker $W_{s}$ and 0 else.
-Furthermore an additional type of decision variables is necessary: $U_{i,j,i^{\prime},j^{\prime}}$, which is set to 1 if Operation $O_{i,j}$ is processed after operation $O_{i^{\prime},j^{\prime}}$ by the same worker.
-To describe the admissible workers for processing operation $O_{i,j}$ on machine $M_{k}$ the index set
-$W_{i,j,k}=\{s:\text{ such that $O_{i,j}$ can be processed on $M_{k}$ by worker$W_{s}$}\}$
-is used.
-4.2.2 Constraints
-Assignment constraints
-These constraints guarantee that all operations $O_{i,j}$ are assigned to exactly one machine and one worker.
-$\sum\nolimits_{k\in M_{i,j}}\sum\nolimits_{s\in W_{i,j,k}}Y_{i,j,k,s}=1\text{ for }1\leq i\leq n,1\leq j\leq n_{i}$
-Sequencing constraints on job level
-$C_{i,1}\geq\sum\nolimits_{k\in M_{i,1}}\sum\nolimits_{s\in W_{i,1,k}}Y_{i,1,k,s}T_{i,1,k,s}\text{ for }1\leq i\leq n$
-$C_{i,j}\geq C_{i,j-1}+\sum\nolimits_{k\in M_{i,j}}\sum\nolimits_{s\in W_{i,j,k}}Y_{i,j,k,s}T_{i,j,k,s}\text{ for }1\leq i\leq n,1<j\leq n_{i}$
-Sequencing constraints on machine level
-$C_{i,j}\geq C_{i^{\prime},j^{\prime}}+T_{i,j,k,s}-L(3-X_{i,j,i^{\prime},j^{\prime}}-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k,s^{\prime}})$
-$C_{i^{\prime},j^{\prime}}\geq C_{i,j}+T_{i^{\prime},j^{\prime},k,s^{\prime}}-L(X_{i,j,i^{\prime},j^{\prime}}+2-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k,s^{\prime}})$
-Both constraints have to hold for operations $O_{i,j}$ with $1\leq i<n$ and $1\leq j\leq n_{i}$ and operation $O_{i^{\prime},j^{\prime}}$ with $i^{\prime}>i$ and $1\leq j^{\prime}\leq n_{i^{\prime}}$, for machines $M_{k}$ with $k\in M_{i,j}\cap M_{i^{\prime},j^{\prime}}$ and for workers $W_{s}$ ($s\in W_{i,j,k}$) and $W_{s^{\prime}}$ ($s^{\prime}\in W_{i^{\prime},j^{\prime},k}$).
-Sequencing constraints on worker level
-$C_{i,j}\geq C_{i^{\prime},j^{\prime}}+T_{i,j,k,s}-L(3-U_{i,j,i^{\prime},j^{\prime}}-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k^{\prime},s})$
-$C_{i^{\prime},j^{\prime}}\geq C_{i,j}+T_{i^{\prime},j^{\prime},k^{\prime},s}-L(U_{i,j,i^{\prime},j^{\prime}}+2-Y_{i,j,k,s}-Y_{i^{\prime},j^{\prime},k^{\prime},s})$
-As in 4.2.2 these two sets of constraints have to hold for operations $O_{i,j}$ with $1\leq i<n$ and $1\leq j\leq n_{i}$ and operation $O_{i^{\prime},j^{\prime}}$ with $i^{\prime}>i$ and $1\leq j^{\prime}\leq n_{i^{\prime}}$.
-But contrary to 4.2.2 they have to hold for machines $M_{k}$ ($k\in M_{i,j}$) and $M_{k^{\prime}}$ ($k^{\prime}\in M_{i^{\prime},j^{\prime}}$) and for workers $W_{s}$ with $s\in W_{i,j,k}\cap W_{i^{\prime},j^{\prime},k^{\prime}}$.
-For the large number $L$, one can use
-$L=\sum\nolimits_{i=1}^{n}\sum\nolimits_{j=1}^{n_{i}}\max\nolimits_{1\leq k\leq m;1\leq s\leq w}(T_{i,j,k,s})$
-The idea of these constraints is the same as in Eq. (9) and (10) explained there in detail. Finally makespan and variable type constraint are the same as in 3.1.
-4.3 Constraint Programming Formulation
-The same as for the MILP formulation holds also for the CP formulation: The FJSSP-W is an obvious generalization of the FJSSP.
-In the following the formulation for CPO solver is presented.
-The necessary changes for using the ORT solver are obvious when the two different formulations for the FJSSP are kept in mind.
-Again, each operation is represented by an interval variable $I(i,j)$ with starting time $start(i,j)$, ending time $end(i,j)$ and duration $duration(i,j)$.
-For the FJSSP-W an alternative interval variable $a\_I(i,j,k,s)$ has to be defined for each admissible machine-worker combination.
-The assignment of an operation $O_{j,k}$ to exactly one machine-worker combination is done by linking the interval variable $I(j,k)$ to an alternative interval variable $a\_I(i,j,k,s)$.
-This can be done by the $alternative$ constraints
-$alternative\left(I(i,j),[a\_I(i,j,k,s):\mbox{ for }k\in M_{i,j},s\in W_{i,j,k}]\right)$
-for $1\leq i\leq n$ and $1\leq j\leq n_{i}$.
-The sequencing constraints on job level are the same as in 3.2:
-$end\_before\_start(I(i,j-1),I(i,j))\mbox{ for }1\leq i\leq n\mbox{ and }2\leq j\leq n_{i}$
-and the sequencing constraints on machine level are very similar to 3.2. For $1\leq k\leq m$ the following constraints are added:
-$no\_overlap\left(a\_I(i,j,k,s):\mbox{ for }1\leq i\leq n,1\leq j\leq n_{i}\mbox{ if }k\in M_{i,j}\mbox{ and }s\in W_{i,j,k}\right)$
-Additionally the sequencing on worker level has to be guaranteed by using another set of $no\_overlap$ constraints. For $1\leq s\leq w$:
-$no\_overlap\left(a\_I(i,j,k,s):\mbox{ for }1\leq i\leq n,1\leq j\leq n_{i},k\in M_{i,j}\mbox{ if }s\in W_{i,j,k}\right)$
-4.4 Simulation-Based Optimization Attempt
-The extension to the encoding for simulation-based approaches stays the same for sequence and machine assignments as in section 3.3.
+C.1 Extension of SBO Encoding for FJSSP-W
+The extension to the encoding for simulation-based approaches stays the same for sequence and machine assignments as in section LABEL:sec:fjssp-sbo.
 To add the worker flexibility, an additional vector is added.
 The vector has the same properties as the machine assignment vector, in that it is in the same fixed order and contains the IDs of the workers assigned to the operation-machine combination chosen to do the task.
-Extending the example used in section 3.3 with 3 available workers for the system, a problem instance could look like Listing 4.
+Extending the example used in section LABEL:sec:fjssp-sbo with 3 available workers for the system, a problem instance could look like Listing 2.
 1s = [1, 2, 1, 1, 2]
@@
 3w = [3, 1, 1, 2, 3]
-Listing 4: Example sequence, assignment and worker vector
+Listing 2: Example sequence, assignment and worker vector
 The resulting schedule looks as depicted in Fig. 11.
 Figure 11: Example schedule produced by Listing 1.
-5 Description of the Benchmarking Environment
-5.1 Implementation details
-The benchmarking environment is implemented in Python [21].
-Several APIs are offered by the environment, including the evaluation of FJSSP and FJSSP-W solutions and their comparison to the known results.
-Additionally, the benchmarking environment contains the best known values for the FJSSP and FJSSP-W instances created with the same parameters than the instances used in this paper.
-The algorithm which turns FJSSP instances into FJSSP-W instances is also provided.
-5.2 Performance evaluation
-For comparisons to other solvers the benchmark environment offers different visualizations of the results over all used benchmark instances.
-The data of the evaluated solvers are included and best known results can be updated over time.
-(a)Refer to caption (b)Refer to caption
-Figure 6: An example plot using example data displaying the portion of instances that lie within a certain gap from the best found solution.
-A gap of 1.0 corresponds to 100% distance.
-In (a), all of the data is displayed, while (b) shows an example with a more zoomed in perspective on the same data (limit on the x-axis as parameter to the visualizer code)
-The example plot shown in Fig.
-6(a) compares the relative deviation (gap) $\delta_{rel}$ (Eq.
-(23)) of the best solution of a solver $C_{fb}=C(\mathbf{y}_{fb})$ for each algorithm from the best-known result $C_{best}$.
-On the y-axis, the plot shows the portion of instances where the solutions were within the given gap to $C_{best}$.
-$\begin{split}\delta_{rel}=\frac{C_{fb}-C_{best}}{C_{best}}.\end{split}$ (23)
-Fig. 6(b) shows the same (on example data) with a zoomed in perspective (limit on the x-axis).
-(a)Refer to caption (b)Refer to caption
-Figure 7: Based on example data, the portion of problem instances per solver that were optimized within 10% of the best found result.
-(a) shows the results in relation to the best solution of all solvers within, while (b) shows the time needed for the instances to reach the threshold in relation to their own best solution.
-The plots shown in Fig.
-7 show the portion of instances which reached the given criteria in $n$ seconds.
-The example plots use the fitness as example, but the same plots can also be done for other criteria, e.g.
-the lower bounds found by a solver.
-The maximal distance to the target value can be given to the visualizer as a parameter.
-In the case of Fig.
-7(a), the plot shows the portion of instances capable of getting within a range of no more than 10% distance from the overall best known result $C_{best}$ within the given time frames.
-Fig.
-7(b) however shows how long it took each solver to get within the range of 10% of it’s own best solution $C_{fb}$ for which portion of the test instances.
-Note that while for Fig.
-6(a) all solvers reach 100% eventually, granted that all instances included in the comparison have been solved by each solver, in Fig.
-7(a) there may be many instances that never reached the threshold.
-The visualizer is implemented in a way that it can handle missing data for the case of Fig.
-6(a).
-The corresponding line for the solver with the missing results would not reach 100%, which is the only way this can happen for these plots (without a limit imposed on the x-axis).
-6 Baseline Experiments
-6.1 Experimental Settings
-For the FJSSP-W, the solvers were tasked with solving all 402 benchmarks.
-The hardware used for the experiments is a Intel Core i7-6700 (3.40GHz) CPU with 16 GB RAM.
-All solvers were given a maximum of 1200 seconds (20 minutes) for each benchmark instance.
-The benchmark instances, including all results are available online [22].
-6.2 Applied Solvers
-All used solvers, including the self implemented GA, were not manually tuned or configured to accommodate any of the benchmark instances and were run in their default settings.
-However, some of the solvers do automated pre-processing and parameterizing on their own, e.g.
-Gurobi or the self implemented GA, which determines some of its parameters automatically based on the benchmark characteristics.
-Hexaly
-Hexaly is a new kind of global optimization solver.
-It offers nonlinear and set-oriented modeling APIs.
-Relying on innovative proprietary algorithms, it is faster and more scalable than Mixed-Integer Programming, Constraint Programming, and Nonlinear Programming solvers.
-[23]
-OR-Tools
-OR-Tools is an open-source software suite for optimization [24] developed by Google.
-It offers several ways to solve constraint programming and linear programming problems.
-The solver used for the results in this paper is the CP-SAT solver, which is a constraint programming solver that uses SAT (satisfiability) methods.
-ILOG CPLEX
-CPLEX by IBM is an optimization software package used to solve several optimization problem types, including integer programming, linear programming, quadratic programming problems, and constraint programming problems.
-CPLEX is used for both its constraint programming capabilities (CPLEX-CP) and its MILP solver capabilities (CPLEX-LP).
-[25]
-Gurobi
-Gurobi is a commercial mathematical solver commonly used to solve MILP problems and is developed by Gurobi Optimization LLC. [26]
-Genetic Algorithm
-To solve the FJSSP-W, an extended version of the GA implemented in [12] is used.
-Additionally, the implementation is done in the programming language C# rather than Python.
-To evaluate the GA properly, all instance are repeated 18 times to account for the randomness of the GA.
-The 18 repetitions are a result of the employed hardware being able to handle 6 instances of the GA running in parallel processes.
-With a maximum running time of 20 minutes, all repetitions for all instances were expected to take 402 hours (16.75 days).
-However, this runtime was reduced by preemptively stopping the GA once it reached the optimal result.
-However, other than the CP and MILP solvers, the GA has no built-in method to determine if a found solution is optimal.
-Therefore, the GA was run as the last solver and relied on results reported by the other solvers as optimal results, if any optimal result was found for a problem instance.
-This helps illustrate which ranges of the different characteristics are easiest to handle.
-For example, it can be observed that the difficulty of a problem increases with a higher number of operations, as well as with a lower duration variety in the processing times.
-6.3 Baseline Results
-Figure 8: CDF plot showing selected characteristics of all problem instances
-The tables for each solver show the number of instances in which the solver reached the best result, the average characteristics of the instances and the respective standard deviations.
-The rows in the tables separate the instances in which the solver was the only solver to reach the best result (Single Best), the instances in which at least one other solver also reached the best result (Shared Best), an overview of all instances for which the solver found the best result (Combined), which includes the 2 prior rows, and lastly the instances for which the solver reached the optimal solution (Optimal).
-Note that since the GA has no way to confirm if a solution is an optimum, the best results for each instances were collected.
-For each solver the instance counts if it reached the known optimum even if it did not identify it as the optimum, as long as one of the solvers confirms the result to be optimal.
-Fig.
-8 shows the cumulative distribution function over all used benchmark instances on selected problem characteristics.
-The vertical lines mark the average characteristic for instances where the best result was achieved by one solver (One Best), by a subset of the solvers (Multiple Best), and instances where all solvers could find the same best result (not necessarily the optimal result).
-Table 2: Characteristics of all benchmarks.
-mean std max min
-$\beta$ 0.2806 0.1749 1.0 0.0667
-$dv$ 0.1496 0.2241 0.95 5.6669e-05
-N 150.1144 111.2573 500 4
-$m$ 13.5124 13.1667 60 2
-6.3.1 MILP
-Gurobi
-The characteristics of the instances Gurobi found the best solutions for are shown in Tab.3.
-Gurobi could not solve any instance that could not also be solved by at least one other solver.
-The instances that can be solved by Gurobi all reach the known optimum (11).
-The instances in question are small in terms of number of machines and number of operations.
-The flexibility is high in comparison to the average flexibility in all benchmarks (Tab.
-2).
-The same is true for the duration variety.
-Table 3: Characteristics of benchmarks in which Gurobi achieved the best result.
-Average Benchmark Attributes
-count $m$ N $\beta$ $dv$
-mean std mean std mean std mean std
-Single Best 0 - - - - - - - -
-Shared Best 11 5.63 2.203 22.63 19.69 0.32 0.14 0.4486 0.3507
-Combined 11 5.63 2.203 22.63 19.69 0.32 0.14 0.4486 0.3507
-Optimal 11 5.63 2.203 22.63 19.69 0.32 0.14 0.4486 0.3507
-CPLEX-LP
-The characteristics of the instances the LP solver included in CPLEX found the best solutions for are shown in Tab.
-4.
-The attributes are similar to the instances solved by Gurobi.
-While CPLEX-LP also could not solve any instances as the only solver, it managed to solve 11 instances to the known optimum.
-The instances solved by CPLEX-LP are smaller than the ones solved by Gurobi on average.
-The flexibility is slightly lower, while the duration variety is significantly higher.
-Table 4: Characteristics of benchmarks in which the LP solver included in CPLEX achieved the best result.
-Average Benchmark Attributes
-count $m$ N $\beta$ $dv$
-mean std mean std mean std mean std
-Single Best 0 - - - - - - - -
-Shared Best 11 4.18 1.83 15.73 14.21 0.32 0.21 0.6506 0.2539
-Combined 11 4.18 1.83 15.73 14.21 0.32 0.21 0.6506 0.2539
-Optimal 11 4.18 1.83 15.73 14.21 0.32 0.21 0.6506 0.2539
-6.3.2 CP
-OR-Tools
-The characteristics of the instances OR-Tools found the best solutions for are shown in Tab.
-5.
-OR-Tools managed to reach the best solution in 15 instances that could not be solved as well by any other solver.
-It also reached the best result for 56 instances for which at least one other solver reached the best result.
-Of the 71 best results reached, 48 are confirmed to be optimal.
-The average amount of machines is slightly higher than for both MLIP approaches, while the standard deviation is similar.
-Compared to CPLEX-CP and the GA the amount and spread of machines is low, but similar to Hexaly.
-For the amount of operations, OR-Tools solves bigger instances than both MILP approaches, but smaller than Hexaly, CPLEX-CP and the GA.
-The flexibility is smaller than the MILP approaches, CPLEX-CP, and GA, and similar to Hexaly.
-In comparison to Hexaly, CPLEX-CP and the GA, OR-Tools requires a higher duration variety to succeed, but can handle a lower duration variety than both MILP approaches, considering that the duration variety does not significantly change for the different categories shown in the table.
-Table 5: Characteristics of benchmarks in which OR-Tools achieved the best result.
-Average Benchmark Attributes
-count $m$ N $\beta$ $dv$
-mean std mean std mean std mean std
-Single Best 15 7.47 2.45 78.0 29.33 0.11 0.06 0.3061 0.2607
-Shared Best 56 6.11 2.96 53.02 40.21 0.39 0.25 0.3912 0.302
-Combined 71 6.39 2.90 58.29 39.35 0.36 0.24 0.3732 0.294
-Optimal 48 5.67 2.23 47.31 33.58 0.43 0.26 0.3960 0.296
-CPLEX-CP
-The characteristics of the instances CPLEX found the best solutions for are shown in Tab.
-6.
-CPLEX-CP managed to find the best solution in 283 cases where no other solver found the best solution.
-The best result was also found for 66 additional benchmark instances for which other solvers also found the best solution.
-69 of the overall 349 found best results are the known optimum of that benchmark instance.
-CPLEX-CP can handle lower duration varieties than the MILP approaches and OR-Tools, as well as slightly lower duration varieties than Hexaly.
-However, it can not handle the lowest duration varieties like the GA.
-In terms of the flexibility, the average flexibility of the instances only CPLEX-CP could solve best is slightly lower than the flexibility where the GA performs best, but higher than for all the other solvers.
-The average values for the instances that only CPLEX-CP solved best are very close to the average instance attributes (Tab.
-2) of all benchmark instances.
-CPLEX-CP can handle a high operation amount in comparison to all other solvers except the GA.
-The average amount of machines is also higher than for the other solvers, the exception again being the GA.
-Table 6: Characteristics of benchmarks in which CPLEX achieved the best result.
-Average Benchmark Attributes
-count $m$ N $\beta$ $dv$
-mean std mean std mean std mean std
-Single Best 283 14.14 12.55 175.96 112.49 0.28 0.15 0.1026 0.177
-Shared Best 55 15.99 21.27 63.53 51.41 0.04 0.06 0.3424 0.294
-Combined 338 14.20 13.89 158.14 114.98 0.014 0.03 0.1361 0.216
-Optimal 69 20.77 21.83 73.99 62.06 0.039 0.05 0.3059 0.303
-6.3.3 SBO
-The characteristics of the instances the GA found the best solutions for are shown in Tab.
-7.
-The GA could find the best solution for 31 instances, 17 of which are the known optimum.
-Any of the 7 best solutions that only the GA could find could also be the optimum, but the GA is lacking any method of confirming the optimum.
-Note that the stable value of 60 for amount of machines point at the Behnke instances [5], which contain the largest problem instances.
-For the instances that only the GA could solve best, the amount of machines and operations are high on average with a big range especially for the amount of operations.
-Paired with the flexibility values, which are slightly higher than CPLEX-CP and significantly higher than for the other solvers, this indicates that the GA is well suited for large and complex instances.
-Additionally, the GA performed better than all other solvers for instances with a low duration variety, which makes these instances harder to solve.
-Through the average metrics for the Shared Best instances, it can be seen that the GA can still also perform well on instances with high duration variety, and higher flexibility.
-Measuring the stability of the GA for every benchmark instance using Eq.
-(24), the average spread around the average result over all instances is 6.35% (median 5.01%).
-$\begin{split}\delta=\cfrac{f_{worst}-f_{best}}{f_{average}}\end{split}$ (24)
-For the instances for which the GA could find the best result overall, the average amount of machines stands out compared to all other solvers.
-The average amount of operations is also higher than the other solvers, though Hexaly comes close to a similar problem size, but with a much lower standard deviation.
-The average flexibility is below OR-Tools but above CPLEX-CP.
-Table 7: Characteristics of benchmarks in which the GA achieved the best result.
-Average Benchmark Attributes
-count $m$ N $\beta$ $dv$
-mean std mean std mean std mean std
-Single Best 7 60.00 0.00 214.29 207.59 0.16 0.02 0.0003 0.0003
-Shared Best 24 11.75 16.73 44.33 71.21 0.34 0.21 0.4393 0.358
-Combined 31 22.65 25.20 82.71 133.13 0.29 0.19 0.3402 0.3648
-Optimal 17 4.65 2.42 12.12 7.12 0.39 0.19 0.6181 0.2603
-The characteristics of the instances for which Hexaly found the best solutions are shown in Tab.
-8.
-Overall, Hexaly found the best solution for 71 of the instances, 37 of which were identified as the optimum (by any of the solvers).
-29 of the best results were on instances where only Hexaly could find the best result, and 42 of the best results were also found by at least one other solver.
-Hexaly performs well on instances with a higher duration variety than the GA and CPLEX-CP but handles a lower $dv$ better than OR-Tools and the MILP approaches.
-Hexaly seems to deal best with the lowest flexibilities out of all solvers.
-The problem size is narrower spread than for the Single Best results of the GA and CPLEX-CP and on average lower, but higher than the other solvers.
-The same applies to the amount of machines.
-Table 8: Characteristics of benchmarks in which Hexaly achieved the best result.
-Average Benchmark Attributes
-count $m$ N $\beta$ $dv$
-mean std mean std mean std mean std
-Single Best 29 10.10 1.92 116.76 52.25 0.07 0.05 0.1424 0.153
-Shared Best 42 6.12 3.19 53.12 52.19 0.24 0.19 0.3887 0.2943
-Combined 71 7.75 3.36 79.11 60.66 0.17 0.17 0.2881 0.2738
-Optimal 37 5.43 2.15 50.27 54.41 0.26 0.19 0.3984 0.289
-6.4 Comparison
-Figure 9: The portion of problem instances per solver that are within a relative gap (Eq. (23)) to the best known result.
-Optimal Feasible Infeasible
-CPLEX_CP 69 331 2
-CLPLEX_LP 11 310 81
-Gurobi 11 305 86
-Hexaly 37 365 0
-OR-Tools 48 349 5
-GA_Best 17 385 0
-GA_Average 14 388 0
-Table 9: Caption
-The results in Fig.
-9 show the relative gap of the results of each solver to the best result found between all solvers.
-While the results in [12] showed that MILP solvers perform better on the FJSSP instances, the results indicated that the GA would start to outperform the MILP solver on more complex problems.
-This can be shown through the data collected using the FJSSP-W instances.
-Considering the best results of the GA as well as the average results of the GA perform better than both used MILP solvers for the FJSSP-W.
-This shows that the GA can scale to more complex scenarios better.
-Fig.
-9 shows that except for the MILP approaches, all solvers are less than 25% worse than the best overall found result on over 80% of all instances.
-Given even more complex scenarios, like the FJSSP-W with added uncertainties for robust optimization, the GA can be expected to increase it’s advantage over the other solvers even further.
-As seen in the results (see Sec.
-6.3), all solvers show tendencies towards a certain category of problem instance.
-While the GA seems to scale best to large, complex problems, the CP solvers are closest to solving the average benchmark instance for the given benchmark collection.
-The known optima found by the solvers all have a higher than average duration variety, showing that a high duration variety indicates instances which are easier to solve across all solvers.
-The same seems to be true for the flexibility.
-Considering the combination of high duration variety and high flexibility, this means that for those instances the assignment problem can be solved easier, leaving the main pressure on solving the sequencing problem.
-However, the average amount of machines and operations on instances that are solved optimally is lower than the average benchmark instance, meaning that mostly small instances with high duration variety are the ones that can be solved by any of the solvers easily.
-Table 10: Average characteristics of optimally solved benchmarks.
-mean std
-$\beta$ 0.4145 0.2251
-$dv$ 0.03872 0.0542
-N 73.9855 62.0568
-$m$ 7.1739 3.3429
-$d_{mean}$ 100 131.3469
-$d_{stdv}$ 52.9665 79.3247
-$d_{span}$ 189 272.8644
-Figure 10: Benchmark attributes of all instances normalized over the maximum value.
-The characteristics of the benchmarks that could be solved optimally are shown in Tab.
-10.
-From this data it can also be seen that the solvers generally seem to profit from an easier assignment problem, with a low number of machines on average and a mid-range flexibility.
-For these smaller instances, a smaller duration variety also appears to be less of an issue, for the same reason.
-The distribution of actual duration values was also considered (Fig.
-10) , since some problem instances deal with duration values smaller than 10 while other instances have duration values up to over 1000, with he maximum span present in the benchmark set being over 1600.
-However, due to the large variation of duration value metrics in the optimally solved instances (Tab.
-10), the actual values do not seem to impact the solvers, as long as the duration variety decreases the number of solutions with the same fitness values and the amount of mutations applied to solutions that do not lead to change in their fitness value.
-In comparison to the average characteristics of all benchmarks (Tab.
-2), the flexibility of the optimally solved instances is higher and the problem size is lower.
-Overall the data implies that the assignment part of the optimization (machines for FJSSP, machines and workers for FJSSP-W) seems more problematic for all solvers but the GA, which additionally explains why the GA scales up better to the FJSSP-W than other solvers.
-For further investigation, it could prove useful to craft a better balanced benchmark collection, varying in problem size (machines, workers, operations) as well as their complexity (flexibility, duration variety), to cover the metrics more evenly for testing and comparison purposes.
-Many of the different benchmark sources also contain multiple instances that cover very similar problems, which results in a lot of experiment runtime is used to test the solvers on the essentially the same problem.
-(a)Refer to caption (b)Refer to caption (c)Refer to caption
-Figure 11: Best solver for benchmark characteristics
-(a)Refer to caption (b)Refer to caption (c)Refer to caption
-Figure 12: Optimal solution found by the different solvers
-(a)Refer to caption (b)Refer to caption (c)Refer to caption
-Figure 13: Best solution based on solver category.
-While MILP approaches could not achieve the single best result for any instance, they could optimize 11 instances the same as all other solvers.
-CP approaches achieved by far the most singled out best results (315), followed by Hexaly (29), then the GA (8).
-39 instances were solved best by a mixture of solver categories and 11 instances were solved the same by all solver categories.
-7 Conclusion
-While the GA only outperforms the MILP approaches on the FJSSP on specific instances  [12], the MILP approaches are not capable of scaling up to the FJSSP-W as well as the GA.
-Except for CPLEX-CP, the same is true for the other approaches, like the CP approach using OR-Tools and Hexaly.
-While both can still keep up with the GA on certain instances, the GA shows better potential in scaling up to more complex scenarios.
-In direct comparison between the CP approach and the GA approach, the characteristics of the benchmark problems solved best by either show some significant differences.
-Namely, the GA deals better with a more difficult assignment problem, with both more machine and workers as well as more operations.
-However, for simpler assignment problems (low $dv$), the CP approach outperforms the GA in runtime.
-While it is likely that the efficiency of the implementation plays a role for those problems, the CP approach does tend to find its final solution earlier on in the optimization process than the GA, which steadily increases its best solution over time.
-This shows that both approaches have an advantage in different scenarios, which can possibly be determined by the problem characteristics, and could be exploited by a hybrid model.
-The capability of the GA to handle low $dv$ values in large problem instances demonstrates its suitability for even more complex use-cases, including FJSSP and FJSSP-W with uncertainties.
-8 Outlook
-To make the current system more applicable to the real-world, the developed solvers need to be more robust towards uncertainties.
-Given that the GA shows stability when scaled to more complex problems, it is well geared to include uncertainty in its optimization process, while other approaches like MILP and CP struggle with uncertainty.
-However, given the efficiency of CP to find good solutions, a hybrid method combining the advantages of both CP and GAs will be considered as well.
-Different types of uncertainties will be included in the future work, like unexpected events (e.g.
-machine breakdowns, worker unavailability) and inaccurate information (e.g.
-stochastic processing times).